### v1

# Generalized Heisenberg Dynamics Revisited

Abstract
Taking as a model the fact that Heisenberg’s matrix mechanics was derived from Hamiltonian mechanics using the correspondence principle, we explore a class of dynamical systems involving discrete variables, with Nambu mechanics as the starting point. Specifically, we reconstruct an extended version of matrix mechanics that describes dynamical systems possessing physical quantities expressed through generalized matrices. Furthermore, we reconfirm that a multiple commutator involving generalized matrices can serve as a discrete (quantized) version of the Nambu bracket or the Jacobian.
1 Introduction
A hundred years ago, Heisenberg formulated matrix mechanics [1], one of the foundational frameworks of quantum mechanics. In its formulation, Bohr’s correspondence principle played a guiding role. The correspondence principle states that, in a certain limit, physical quantities in quantum theory correspond to those in classical theory. In matrix mechanics (Heisenberg dynamics), physical quantities are represented by Hermitian matrices [2, 3], and the theoretical structure exhibits similarities to Hamiltonian mechanics (Hamiltonian dynamics). For example, Heisenberg’s equation of motion, which is the fundamental equation of matrix mechanics, takes the form of Hamilton’s canonical equations with the Poisson brackets replaced by commutators (divided by the imaginary unit and the reduced Planck constant). The existence of such similarities is considered to be one of the reasons why the correspondence principle functions effectively.
A few months after the advent of matrix mechanics, Schrödinger proposed wave mechanics [4, 5, 6, 7], which is based on the Schrödinger equation as its fundamental equation, and successfully reproduced the results of matrix mechanics. Subsequently, London, Jordan and Dirac demonstrated the equivalence of these two formulations through transformation theory [8, 9, 10]. At present, within the framework of quantum mechanics, matrix mechanics is understood as the theoretical formulation in the Heisenberg picture, in which the operator associated with energy (the Hamiltonian) is diagonalized, while wave mechanics is understood as the formulation in the Schrödinger picture, in which the operator associated with position is diagonalized.
Today, quantum mechanics and its extension to field theory, known as quantum field theory (collectively referred to here as quantum theory), are widely utilized across various areas of physics and have achieved remarkable success. In light of such success, there naturally arises an interest in exploring the foundations of quantum theory, investigating the scope of its applicability, and examining the potential for its extension. Indeed, motivated by such considerations, attempts have been made to extend matrix mechanics, and its fundamental structure has been studied from an algebraic perspective [11, 12, 13, 14]. This generalized mechanics is regarded as an extension of matrix mechanics, as it constructs physical quantities using an extended version of matrices, starting from generalized relations inspired by Bohr’s frequency condition and Ritz’s combination rule (see Figure 1).
Figure 1: Extensions of Hamiltonian mechanics and matrix mechanics.
Furthermore, this generalized matrix mechanics exhibits similarities to Nambu mechanics (Nambu dynamics) and can be regarded as a quantum version of Nambu mechanics. Here, Nambu mechanics is a generalization of Hamiltonian dynamics through the extension of phase space based on the Liouville theorem [15] (see Figure 1). Specifically, in Ref. [13], the relationship between Nambu mechanics based on canonical triplets and the dynamics involving extended matrices with three indices is discussed from the viewpoint of the correspondence principle. However, this analysis is somewhat ad hoc and remains insufficient. Moreover, it is unclear whether a similar correspondence holds in the general case of dynamics involving extended matrices with $n$ indices. In Ref. [14], the theoretical formulation of dynamics based on extended matrices with $n$ indices was explored, but explicit expressions of frequency-related equations in terms of the eigenvalues of the Hamiltonians were not provided. In other words, further clarification of the theoretical structure of dynamics based on extended matrices with $n$ indices remains an open problem.
In this study, we address the question: “What kind of theoretical structure emerges when the correspondence principle is applied to Nambu mechanics?”. Answering this question is of considerable physical significance. This is because such an investigation can shed light on the ‘quantum’ or discrete version of the Nambu bracket, namely the quantum analogue of the Jacobian, potentially leading to a formulation of quantum theory for extended objects222 See Refs. [16, 17] for a comprehensive review of some aspects of Nambu mechanics and its applications in string/M-theory.. Furthermore, obtaining more concrete expressions for frequencies is expected to provide a clearer understanding of the underlying theoretical structure of an extended mechanics.
In this paper, we construct a framework of mechanics containing discrete variables, using the correspondence principle as the guiding principle and taking Nambu mechanics as the starting point. Specifically, we reconstruct an extended version of Heisenberg’s matrix mechanics that describes dynamical systems whose physical quantities are represented by generalized matrices. As a byproduct, we reconfirm that a multiple commutator involving generalized matrices can serve as the quantum or discrete analogue of the Nambu bracket, namely the quantum version of Jacobian.
The outline of this paper is as follows. In the next section, we review the structure of Hamiltonian dynamics and demonstrate how matrix mechanics can be derived from it using the correspondence principle. In section 3, we construct a dynamical framework involving extended matrices with three indices, starting from Nambu mechanics based on canonical triplets and applying the correspondence principle. In section 4, we develop a dynamical theory involving extended matrices with $n$ indices and study the properties of $n$-fold commutators defined on generalized matrices. In the final section, we present our conclusions and discussions. A detailed derivation of the equations presented in section 4 is provided in the Appendix A.
2 From Hamiltonian dynamics to Heisenberg dynamics
2.1 Hamiltonian dynamics
First, we review the structure of Hamiltonian dynamics [18], using a periodic motion described by the canonical variables $q=q(t)$ and $p=p(t)$. These variables satisfy Hamilton’s canonical equations of motion:
$\frac{dq}{dt}=\frac{\partial H}{\partial p},~{}~{}\frac{dp}{dt}=-\frac{\partial H}{\partial q},$ (1)
where $H=H(q,p)$ is the Hamiltonian. Using eq. (1), we find that a general variable $A=A(q,p)$ satisfies the equation:
$\frac{dA}{dt}=\{A,H\}_{\rm PB},$ (2)
where $\{A,H\}_{\rm PB}$ is the Poisson bracket defined by
$\{A,H\}_{\rm PB}\equiv\frac{\partial(A,H)}{\partial(q,p)}=\frac{\partial A}{\partial q}\frac{\partial H}{\partial p}-\frac{\partial H}{\partial q}\frac{\partial A}{\partial p}.$ (3)
From eqs. (2) and (3), we see that $H$ conserves, i.e., $dH/dt=\{H,H\}_{\rm PB}=0$.
The infinitesimal canonical transformation is given by
$\delta_{G}A=\varepsilon\{A,G\}_{\rm PB},$ (4)
where $\varepsilon$ is an infinitesimal constant and $G=G(q,p)$ is a generating function. Then, using the Jacobi identity:
$\{\{A,G\}_{\rm PB},H\}_{\rm PB}+\{\{G,H\}_{\rm PB},A\}_{\rm PB}+\{\{H,A\}_{\rm PB},G\}_{\rm PB}=0,$ (5)
it is shown that the equation is invariant under the canonical transformation such that
$\frac{dA^{\prime}}{dt}=\{A^{\prime},H\}_{\rm PB},$ (6)
where $A^{\prime}=A+\delta_{G}A$. From eqs. (2) and (4), we find that the time evolution of variables is realized by the canonical transformation whose generation function is the Hamiltonian.
Let us carry out the canonical transformation $(q,p)\to(\theta,J)$, in which $\theta$ becomes a cyclic coordinate, using a generating function $W=W(q,J)$. The canonical pairs are related to each other in the following way,
$pdq-Hdt=-\theta dJ-H^{\prime}dt+dW,$ (7)
where $H^{\prime}=H^{\prime}(J)$ is a new Hamiltonian. We notice that $H^{\prime}$ does not depend on $\theta$. New canonical variables $\theta$ and $J$ satisfy Hamilton’s canonical equations of motion:
$\frac{d\theta}{dt}=\frac{\partial H^{\prime}}{\partial J},~{}~{}\frac{dJ}{dt}=-\frac{\partial H^{\prime}}{\partial\theta}=0.$ (8)
From eq. (8), we find that $J$ becomes a constant and then $\theta=\theta(t)$ is given by
$\theta(t)=\nu t+\theta_{0},$ (9)
where $\theta_{0}$ is a constant and $\nu$ is also a constant given by
$\nu=\frac{dE}{dJ}.$ (10)
In eq. (10), $E=E(J)$ is the value of $H^{\prime}$.
From eq. (7), $p$, $\theta$, $H$ and $H^{\prime}$ are given by
$p=\frac{\partial W}{\partial q},~{}~{}\theta=\frac{\partial W}{\partial J},~{}~{}H=H^{\prime}=E,$ (11)
respectively.
[…]
from the relation:
[…]
where we use eq. (11) and C stands for a closed trajectory on the phase space. The variable $\theta$ and $J$ are called the angle variable and the action variable, respectively. Inserting the momentum $p=p(q,E)$ derived from $H(q,p)=E$ into eq. (12), we see that $J$ becomes a function of $E$. Differentiating $J=J(E)$ by $E$ and using $(\partial p/\partial E)(\partial H/\partial p)=1$ and ${{dq}/{dt}={\partial H}/{\partial p}}$, we obtain the relation:
[…]
where $T$ is a period. From eq. (14), eq. (10) is rederived as
$\frac{dE(J)}{dJ}=\frac{1}{T}=\nu.$ (15)
Using eq. (14) and $T=1/\nu$, $J$ is rewritten as
[…]
In the classical electromagnetism, electromagnetic waves are emitted from a rotating charged particle and the frequency of light $\nu_{\rm cl}$ follows the relation:
$\nu_{\rm cl}=\nu\Delta n=\frac{dE(J)}{dJ}\Delta n,$ (17)
where $\nu$ is a number of rotation of the charged particle and $\Delta n$ is a natural number. It is known that the formula (17) does not hold microscopically and it is modified at the quantum level.
2.2 Heisenberg dynamics
Let us construct Heisenberg dynamics from Hamiltonian dynamics, taking the correspondence principle as a guiding principle. To go into more detail about the correspondence principle, it states that results in classical physics should be derived as limiting cases of those in quantum physics. In concretely, there is a one-to-one correspondence between quantum quantities and classical quantities, and values of quantum ones and relations among quantum ones approach values of classical ones and relations among classical ones in the limit where a quantum number $n$ becomes large. This can be expressed as
$A_{n}\xrightarrow{n:{\rm large}}A_{\rm cl},~{}~{}B_{n}\xrightarrow{n:{\rm large}}B_{\rm cl},~{}~{}~{}~{}A_{n}=B_{n}\xrightarrow{n:{\rm large}}A_{\rm cl}=B_{\rm cl},$ (18)
where $A_{n}$ and $B_{n}$ are quantum counterparts of classical quantities $A_{\rm cl}$ and $B_{\rm cl}$, respectively.
We assume that the quantum value of energy $E_{n}$ and the classical one $E(J)$ are related to each other as
$E_{n}=E(hn)$ (19)
through the Bohr-Sommerfeld quantization condition:
[…]
where $h$ is the Planck constant and $n$ is a natural number.
In this case, we obtain the following correspondence relation:
$\frac{E_{n}-E_{n-\Delta n}}{h}=\frac{E(hn)-E(h(n-\Delta n))}{h}=\frac{E(hn)-E(hn-h\Delta n)}{h\Delta n}\Delta n$ (21)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\xrightarrow{n\gg\Delta n}\frac{1}{h}\frac{dE(hn)}{dn}\Delta n=\frac{dE(J)}{dJ}\Delta n=\nu\Delta n=\nu_{\rm cl},$
and hence we realize that $(E_{n}-E_{n-\Delta n})/h$ must be the frequency of photon in the quantum world. In this way, we obtain Bohr’s frequency condition:
$\nu_{m~{}\!\!n}=\frac{E_{m}-E_{n}}{h},$ (22)
where $m$ and $n$ are positive integers. Eq. (22) is the quantum counterpart of eq. (17), and this formula with specific energy eigenvalues explains experimental data of spectrum of light emitting from atoms very well. For example, the energy eigenvalues are given by $E_{n}=-{hcR}/{n^{2}}$ ($R$ : the Rydberg constant, $c$ : the speed of light) for the hydrogen atom.
Next, let us derive an equation of motion for quantum quantities based on the frequency condition (22). We consider a periodic motion with the period $T$ and suppose that quantities have two indices. Then, the quantities $(A_{1})_{m~{}\!\!n}(t)$ and $(A_{2})_{m~{}\!\!n}(t)$ have a feature that $(A_{1})_{m~{}\!\!n}(t+T)=(A_{1})_{m~{}\!\!n}(t)$ and $(A_{2})_{m~{}\!\!n}(t+T)=(A_{2})_{m~{}\!\!n}(t)$ and can be described as
$(A_{1})_{m~{}\!\!n}(t)=(A_{1})_{m~{}\!\!n}(0)e^{2\pi i\nu_{m~{}\!\!n}t},~{}~{}(A_{2})_{m~{}\!\!n}(t)=(A_{2})_{m~{}\!\!n}(0)e^{2\pi i\nu_{m~{}\!\!n}t},$ (23)
where $\nu_{m~{}\!\!n}$ follows the condition (22) and is also written by $\nu_{m~{}\!\!n}=n_{m~{}\!\!n}/T$ ($n_{m~{}\!\!n}$ : an integer).
If we require that the product $(A_{1}A_{2})_{m~{}\!\!n}(t)$ should take the same form as eq. (23), i.e., $(A_{1}A_{2})_{m~{}\!\!n}(t)=(A_{1}A_{2})_{m~{}\!\!n}(0)e^{2\pi i\nu_{m~{}\!\!n}t}$, $(A_{1})_{m~{}\!\!n}(t)$ and $(A_{2})_{m~{}\!\!n}(t)$ become components of matrices whose product is defined by
$(A_{1}A_{2})_{m~{}\!\!n}(t)\equiv\sum_{k}(A_{1})_{m~{}\!\!k}(t)(A_{2})_{k~{}\!\!n}(t),$ (24)
where we use Ritz’s combination rule:
$\nu_{m~{}\!\!n}=\nu_{m~{}\!\!k}+\nu_{k~{}\!\!n}.$ (25)
It is easy to see that the rule (25) holds for $\nu_{m~{}\!\!n}$ given by eq. (22).
Because the diagonal components of $(A_{1})_{m~{}\!\!n}(t)$ and $(A_{2})_{m~{}\!\!n}(t)$ do not depend on time as seen from eqs. (22) and (23), any conserved quantity $C_{m~{}\!\!n}$ satisfying $dC_{m~{}\!\!n}/dt=0$ can be written as $C_{m~{}\!\!n}=c_{m}\delta_{m~{}\!\!n}$ with a constant $c_{m}$. In a closed system, the Hamiltonian conserves, and hence its element $H_{m~{}\!\!n}$ takes the form:
$H_{m~{}\!\!n}=E_{m}\delta_{m~{}\!\!n},$ (26)
using the energy eigenvalue $E_{m}$. In matrix mechanics, the task is to diagonalize the Hamiltonian, which is a function of matrices, in order to determine its eigenvalues.
Using eqs. (22), (23) and (26), we derive the equation:
$\frac{d}{dt}A_{m~{}\!\!n}(t)$ $=2\pi i\nu_{m~{}\!\!n}A_{m~{}\!\!n}(t)=2\pi i\frac{E_{m}-E_{n}}{h}A_{m~{}\!\!n}(t)$
$=\frac{2\pi i}{h}\left(\sum_{k}E_{m}\delta_{m~{}\!\!k}A_{k~{}\!\!n}(t)-\sum_{k}A_{m~{}\!\!k}(t)E_{k}\delta_{k~{}\!\!n}\right)$
$=\frac{2\pi i}{h}\sum_{k}\left(H_{m~{}\!\!k}A_{k~{}\!\!n}(t)-A_{m~{}\!\!k}(t)H_{k~{}\!\!n}\right)$
$=\frac{2\pi i}{h}[H,A(t)]_{m~{}\!\!n}=\frac{1}{i\hbar}[A(t),H]_{m~{}\!\!n},$ (27)
where ${\hbar(\equiv{h}/({2\pi}))}$ is the reduced Planck constant and $[A(t),H]_{m~{}\!\!n}$ is a commutator defined by
$[A(t),H]_{m~{}\!\!n}\equiv\sum_{k}\left(A_{m~{}\!\!k}(t)H_{k~{}\!\!n}-H_{m~{}\!\!k}A_{k~{}\!\!n}(t)\right).$ (28)
In this way, we have arrived at Heisenberg’s equation of motion:
$\frac{d}{dt}A_{m~{}\!\!n}(t)=\frac{1}{i\hbar}[A(t),H]_{m~{}\!\!n},$ (29)
as the quantum counterpart of eq. (2).
The infinitesimal unitary transformation is defined by
$\delta_{G}A_{m~{}\!\!n}=\varepsilon[A,G]_{m~{}\!\!n},$ (30)
where $\varepsilon$ is an infinitesimal constant and $G_{m~{}\!\!n}$ is a generator. Then, using the Jacobi identity:
$\left[[A,G],H\right]_{m~{}\!\!n}+\left[[G,H],A\right]_{m~{}\!\!n}+\left[[H,A],G\right]_{m~{}\!\!n}=0,$ (31)
it is shown that the equation is invariant under the above transformation such that
$\frac{d}{dt}A^{\prime}_{m~{}\!\!n}=[A^{\prime},H]_{m~{}\!\!n},$ (32)
where $A^{\prime}_{m~{}\!\!n}=A_{m~{}\!\!n}+\delta_{G}A_{m~{}\!\!n}$. From eqs. (29) and (30), we find that the time evolution of variables is generated by the unitary transformation whose generator is the Hamiltonian.
2.2.1 Fermionic harmonic oscillator
We study a harmonic oscillator relating to fermionic variables as an example. They are represented by the $2\times 2$ Hermitian matrices whose components are given by
$\xi_{m~{}\!\!n}(t)=\frac{1}{\sqrt{2}}|\varepsilon_{m~{}\!\!n}|e^{-i\omega\varepsilon_{m~{}\!\!n}t},~{}~{}\eta_{m~{}\!\!n}(t)=-\frac{i}{\sqrt{2}}\varepsilon_{m~{}\!\!n}e^{-i\omega\varepsilon_{m~{}\!\!n}t},$ (33)
where $m$ and $n$ run from $1$ to $2$, $\omega$ is the angular frequency, $\varepsilon_{1~{}\!\!2}=-\varepsilon_{2~{}\!\!1}=1$ and $\varepsilon_{1~{}\!\!1}=\varepsilon_{2~{}\!\!2}=0$. These variables satisfy the anti-commutation relations:
$\left\{\xi(t),\xi(t)\right\}_{m~{}\!\!n}=\delta_{m~{}\!\!n},~{}~{}\left\{\eta(t),\eta(t)\right\}_{m~{}\!\!n}=\delta_{m~{}\!\!n},~{}~{}\left\{\xi(t),\eta(t)\right\}_{m~{}\!\!n}=0,$ (34)
where $\{A,B\}_{m~{}\!\!n}\equiv\sum_{k}(A_{m~{}\!\!k}B_{k~{}\!\!n}+B_{m~{}\!\!k}A_{k~{}\!\!n})$.
Fermionic variables become solutions of Heisenberg’s equation of motion:
$\frac{d}{dt}\xi_{m~{}\!\!n}(t)=\frac{1}{i\hbar}[\xi(t),H]_{m~{}\!\!n},~{}~{}\frac{d}{dt}\eta_{m~{}\!\!n}(t)=\frac{1}{i\hbar}[\eta(t),H]_{m~{}\!\!n},$ (35)
where the Hamiltonian is given by
$H_{m~{}\!\!n}=i\hbar\omega\sum_{k=1}^{2}\xi_{m~{}\!\!k}(t)\eta_{k~{}\!\!n}(t)=-\frac{\hbar\omega}{2}\sum_{k=1}^{2}\varepsilon_{m~{}\!\!k}~{}\!\delta_{m~{}\!\!n}.$ (36)
Since $\xi_{m~{}\!\!n}(t)$ and $\eta_{m~{}\!\!n}(t)$ follow the equations:
$\frac{d}{dt}\xi_{m~{}\!\!n}(t)=\omega{\eta}_{m~{}\!\!n}(t),~{}~{}\frac{d}{dt}\eta_{m~{}\!\!n}(t)=-\omega{\xi}_{m~{}\!\!n}(t),$ (37)
they obey the equation for simple harmonic motion:
$\frac{d^{2}}{dt^{2}}\xi_{m~{}\!\!n}(t)=-\omega^{2}{\xi}_{m~{}\!\!n}(t),~{}~{}\frac{d^{2}}{dt^{2}}\eta_{m~{}\!\!n}(t)=-\omega^{2}{\eta}_{m~{}\!\!n}(t).$ (38)
The same system is described by the representation matrices of the ladder operators:
$C_{m~{}\!\!n}(t)\equiv\frac{1}{\sqrt{2}}\left(\xi_{m~{}\!\!n}(t)+i\eta_{m~{}\!\!n}(t)\right)=\frac{1}{2}\left(|\varepsilon_{m~{}\!\!n}|+\varepsilon_{m~{}\!\!n}\right)e^{-i\omega t},~{}~{}$ (39)
$C^{\dagger}_{m~{}\!\!n}(t)\equiv\frac{1}{\sqrt{2}}\left(\xi_{m~{}\!\!n}(t)-i\eta_{m~{}\!\!n}(t)\right)=\frac{1}{2}\left(|\varepsilon_{m~{}\!\!n}|-\varepsilon_{m~{}\!\!n}\right)e^{i\omega t}.$ (40)
These variables obey the anti-commutation relations:
$\{C(t),C^{\dagger}(t)\}_{m~{}\!\!n}=\delta_{m~{}\!\!n},~{}~{}\{C(t),C(t)\}_{m~{}\!\!n}=0,~{}~{}\{C^{\dagger}(t),C^{\dagger}(t)\}_{m~{}\!\!n}=0,$ (41)
and satisfy Heisenberg’s equation of motion:
$\frac{d}{dt}C_{m~{}\!\!n}(t)=\frac{1}{i\hbar}[C(t),H]_{m~{}\!\!n},~{}~{}\frac{d}{dt}C^{\dagger}_{m~{}\!\!n}(t)=\frac{1}{i\hbar}[C^{\dagger}(t),H]_{m~{}\!\!n},$ (42)
where the Hamiltonian is written by
$H_{m~{}\!\!n}=\hbar\omega[C^{\dagger}(t),C(t)]_{m~{}\!\!n}=\hbar\omega\left(\sum_{k=1}^{2}C^{\dagger}_{m~{}\!\!k}(t)C_{k~{}\!\!n}(t)-\frac{1}{2}\delta_{m~{}\!\!n}\right).$ (43)
Since $C_{m~{}\!\!n}(t)$ and $C^{\dagger}_{m~{}\!\!n}(t)$ follow the equations:
$\frac{d}{dt}C_{m~{}\!\!n}(t)=-i\omega C_{m~{}\!\!n}(t),~{}~{}\frac{d}{dt}C^{\dagger}_{m~{}\!\!n}(t)=i\omega C^{\dagger}_{m~{}\!\!n}(t),$ (44)
they also obey the equation for simple harmonic motion.
3 From Nambu dynamics to generalized Heisenberg dynamics
3.1 Nambu dynamics
First, we review the structure of Nambu dynamics [15], using a periodic motion described by the canonical triplet $x=x(t)$, $y=y(t)$ and $z=z(t)$. These variables satisfy a generalization of Hamilton’s canonical equations of motion called the Nambu equation:
$\frac{dx}{dt}=\frac{\partial(H_{1},H_{2})}{\partial(y,z)},~{}~{}\frac{dy}{dt}=\frac{\partial(H_{1},H_{2})}{\partial(z,x)},~{}~{}\frac{dz}{dt}=\frac{\partial(H_{1},H_{2})}{\partial(x,y)},$ (45)
where $H_{1}=H_{1}(x,y,z)$ and $H_{2}=H_{2}(x,y,z)$ are the Hamiltonians. Using eq. (45), we find that a general variable $A=A(x,y,z)$ satisfies the equation:
$\frac{dA}{dt}=\{A,H_{1},H_{2}\}_{\rm NB},$ (46)
where $\{A,H_{1},H_{2}\}_{\rm NB}$ is the Nambu bracket defined by
$\{A,H_{1},H_{2}\}_{\rm NB}$ $\equiv\frac{\partial(A,H_{1},H_{2})}{\partial(x,y,z)}$
$=\frac{\partial A}{\partial x}\frac{\partial H_{1}}{\partial y}\frac{\partial H_{2}}{\partial z}+\frac{\partial H_{1}}{\partial x}\frac{\partial H_{2}}{\partial y}\frac{\partial A}{\partial z}+\frac{\partial H_{2}}{\partial x}\frac{\partial A}{\partial y}\frac{\partial H_{1}}{\partial z}$
$~{}~{}~{}-\frac{\partial H_{2}}{\partial x}\frac{\partial H_{1}}{\partial y}\frac{\partial A}{\partial z}-\frac{\partial H_{1}}{\partial x}\frac{\partial A}{\partial y}\frac{\partial H_{2}}{\partial z}-\frac{\partial A}{\partial x}\frac{\partial H_{2}}{\partial y}\frac{\partial H_{1}}{\partial z}.$ (47)
From eqs. (46) and (47), we see that $H_{1}$ and $H_{2}$ conserve, i.e., $dH_{1}/dt=\{H_{1},H_{1},H_{2}\}_{\rm NB}=0$ and $dH_{2}/dt=\{H_{2},H_{1},H_{2}\}_{\rm NB}=0$.
The infinitesimal generalized canonical transformation is defined by
$\delta_{G}A=\varepsilon\{A,G_{1},G_{2}\}_{\rm NB},$ (48)
where $\varepsilon$ is an infinitesimal constant and $G_{1}=G_{1}(x,y,z)$ and $G_{2}=G_{2}(x,y,z)$ are the generating functions. Then, using the fundamental identity [19]:
$\{\{A,G_{1},G_{2}\}_{\rm NB},H_{1},H_{2}\}_{\rm NB}$ (49)
=$~{}~{}~{}=\{\{A,H_{1},H_{2}\}_{\rm NB},G_{1},G_{2}\}_{\rm NB}+\{A,\{G_{1},H_{1},H_{2}\}_{\rm NB},G_{2}\}_{\rm NB}+\{A,G_{1},\{G_{2},H_{1},H_{2}\}_{\rm NB}\}_{\rm NB},$
it is shown that the equation is invariant under the generalized canonical transformation such that
$\frac{dA^{\prime}}{dt}=\{A^{\prime},H_{1},H_{2}\}_{\rm NB},$ (50)
where $A^{\prime}=A+\delta_{G}A$. From eqs. (46) and (48), we see that the time evolution of variables is realized by the generalized canonical transformation whose generation functions are the Hamiltonians.
Let us carry out the generalized canonical transformation $(x,y,z)\to(\theta,J_{1},J_{2})$, in which $\theta$ becomes a cyclic coordinate, using a function $W_{(1)}=W_{(1)}(y,z,J_{1},J_{2})=W_{1}dW_{2}$. The canonical triplets are related to each other in the following way,
$xdy\wedge dz-H_{1}dH_{2}\wedge dt=\theta dJ_{1}\wedge dJ_{2}-H^{\prime}_{1}dH^{\prime}_{2}\wedge dt+dW_{1}\wedge dW_{2},$ (51)
where $H^{\prime}_{1}=H^{\prime}_{1}(J_{1},J_{2})$ and $H^{\prime}_{2}=H^{\prime}_{2}(J_{1},J_{2})$ are new Hamiltonians. We notice that $H^{\prime}_{1}$ and $H^{\prime}_{2}$ do not depend on $\theta$. New canonical variables $\theta$, $J_{1}$ and $J_{2}$ satisfy the Nambu equation:
$\frac{d\theta}{dt}=\frac{\partial(H^{\prime}_{1},H^{\prime}_{2})}{\partial(J_{1},J_{2})},~{}~{}\frac{dJ_{1}}{dt}=\frac{\partial(H^{\prime}_{1},H^{\prime}_{2})}{\partial(J_{2},\theta)}=0,~{}~{}\frac{dJ_{2}}{dt}=\frac{\partial(H^{\prime}_{1},H^{\prime}_{2})}{\partial(\theta,J_{1})}=0.$ (52)
From eq. (52), we find that $J_{1}$ and $J_{2}$ become constants and then $\theta=\theta(t)$ is given by
$\theta(t)=\nu t+\theta_{0},$ (53)
where $\theta_{0}$ is a constant and $\nu$ is also a constant given by
$\nu=\frac{\partial(E_{1},E_{2})}{\partial(J_{1},J_{2})}.$ (54)
In eq. (54), $E_{1}=E_{1}(J_{1},J_{2})$ and $E_{2}=E_{2}(J_{1},J_{2})$ are the values of $H^{\prime}_{1}$ and $H^{\prime}_{2}$, respectively.
From eq. (51), $x$, $\theta$, $H_{1}$, $H_{2}$, $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are given by
$x=\frac{\partial(W_{1},W_{2})}{\partial(y,z)},~{}~{}\theta=-\frac{\partial(W_{1},W_{2})}{\partial(J_{1},J_{2})},~{}~{}H_{1}dH_{2}=H^{\prime}_{1}dH^{\prime}_{2},$ (55)
respectively. The period $T$ is given by the relation:
[…]
where we use the relation $({\partial(y,z)}/{\partial(E_{1},E_{2})})({\partial(H_{1},H_{2})}/{\partial(y,z)})=1$ and $dH_{1}\wedge dH_{2}=dH^{\prime}_{1}\wedge dH^{\prime}_{2}=dE_{1}\wedge dE_{2}$. When $\nu$ is regarded as a frequency, the following relation is obtained
[…]
using eqs. (54) and (56). Integrating eq. (57) over $E_{1}$ and $E_{2}$, we derive the relation:
[…]
It is natural to impose the following quantization condition on the canonical triplet,
[…]
as a reference of the condition (20) in the old quantum theory.
We study the relationship between Nambu dynamics and Hamiltonian dynamics. Since $H_{1}$ and $H_{2}$ conserve, a particle moves on the path as the intersection of two surfaces described by $H_{1}=E_{1}$ and $H_{2}=E_{2}$. If we carry on the generalized canonical transformation $(x,y,z)\to(x^{\prime},y^{\prime},z^{\prime})$ with $z^{\prime}=H_{2}$, the Nambu equation becomes as
$\frac{dx^{\prime}}{dt}=\frac{\partial H_{1}}{\partial y^{\prime}},~{}~{}\frac{dy^{\prime}}{dt}=-\frac{\partial H_{1}}{\partial x^{\prime}},~{}~{}\frac{dz^{\prime}}{dt}=0,$ (60)
and it turns out to be Hamilton’s equation of motion by identifying $x^{\prime}$, $y^{\prime}$ and $H_{1}$ as $q$, $p$ and $H$, respectively.
3.2 Generalized Heisenberg dynamics
Nambu examined how to quantize Nambu mechanics and found it to be difficult [15]. After that various attempts on the quantization of Nambu mechanics or the Nambu bracket have been carried out [19, 20, 21, 22, 23, 24].
Let us formulate an extension of Heisenberg dynamics from Nambu dynamics, using the correspondence principle (see Figure 2).
Figure 2: Strategy for exploring new dynamics.
We assume that there is a one-to-one correspondence between values of $E_{1}(J_{1},J_{2})$ and $E_{2}(J_{1},J_{2})$ in Nambu dynamics and some discrete values $(E_{1})_{l~{}\!\!m}$ and $(E_{2})_{l~{}\!\!m}$ in its extension, and these dicrete values are anti-symmetric such as
$(E_{1})_{m~{}\!\!l}=-(E_{1})_{l~{}\!\!m},~{}~{}(E_{2})_{m~{}\!\!l}=-(E_{2})_{l~{}\!\!m}$ (61)
and follow the combination rule:
$(E_{1})_{l~{}\!\!l-\Delta l}=(E_{1})_{l~{}\!\!m}+(E_{1})_{m~{}\!\!l-\Delta l}=(E_{1})_{l~{}\!\!m}-(E_{1})_{l-\Delta l~{}\!\!m},~{}~{}$ (62)
$(E_{2})_{m~{}\!\!m-\Delta m}=(E_{2})_{m~{}\!\!l}+(E_{2})_{l~{}\!\!m-\Delta m}=(E_{2})_{m~{}\!\!l}-(E_{2})_{m-\Delta m~{}\!\!l},$ (63)
where $\Delta l$ and $\Delta m$ are natural numbers. In this case, we conjecture the correspondence relation:
$\left(\frac{(E_{1})_{l~{}\!\!l-\Delta l}}{\Delta l}\frac{(E_{2})_{m~{}\!\!m-\Delta m}}{\Delta m}-\frac{(E_{2})_{l~{}\!\!l-\Delta l}}{\Delta l}\frac{(E_{1})_{m~{}\!\!m-\Delta m}}{\Delta m}\right)\Delta l\Delta m$ (64)
$~{}~{}~{}~{}~{}=\left(\frac{(E_{1})_{l~{}\!\!m}-(E_{1})_{l-\Delta l~{}\!\!m}}{\Delta l}\frac{(E_{2})_{m~{}\!\!l}-(E_{2})_{m-\Delta m~{}\!\!l}}{\Delta m}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.-\frac{(E_{2})_{l~{}\!\!m}-(E_{2})_{l-\Delta l~{}\!\!m}}{\Delta l}\frac{(E_{1})_{m~{}\!\!l}-(E_{1})_{m-\Delta m~{}\!\!l}}{\Delta m}\right)\Delta l\Delta m$
$~{}~{}~{}~{}~{}\xrightarrow{l\gg\Delta l,~{}m\gg\Delta m}\frac{\partial(E_{1},E_{2})}{\partial(J_{1},J_{2})}\Delta l\Delta m=\nu\Delta l\Delta m,$
and obtain a frequency condition such as
$\nu^{(0)}_{l~{}\!\!m~{}\!\!n}=\beta\left\{(E_{1})_{l~{}\!\!n}(E_{2})_{m~{}\!\!n}-(E_{2})_{l~{}\!\!n}(E_{1})_{m~{}\!\!n}\right\},$ (65)
where $\beta$ is a constant and we choose $n=l-\Delta l=m-\Delta m$. Eq. (65) is regarded as a counterpart of the relation (54).
Using eqs. (61), (62), (63) and (65), we derive the relation:
$\nu^{(0)}_{l~{}\!\!m~{}\!\!n}=\nu^{(0)}_{l~{}\!\!m~{}\!\!k}+\nu^{(0)}_{m~{}\!\!n~{}\!\!k}+\nu^{(0)}_{n~{}\!\!l~{}\!\!k},$ (66)
which represents that $\nu^{(0)}_{l~{}\!\!m~{}\!\!n}$ is a 3-cocycle, i.e., $(\delta\nu^{(0)})_{l~{}\!\!m~{}\!\!n~{}\!\!k}=0$. Here, $\delta$ is a coboundary operator defined by
$(\delta\nu^{(0)})_{l~{}\!\!m~{}\!\!n~{}\!\!k}\equiv\nu^{(0)}_{m~{}\!\!n~{}\!\!k}-\nu^{(0)}_{l~{}\!\!n~{}\!\!k}+\nu^{(0)}_{l~{}\!\!m~{}\!\!k}-\nu^{(0)}_{l~{}\!\!m~{}\!\!n}.$ (67)
From eqs. (65) and (66), we find that $\nu^{(0)}_{l~{}\!\!m~{}\!\!n}$ is totally anti-symmetric, i.e., $\nu^{(0)}_{l~{}\!\!m~{}\!\!n}=\nu^{(0)}_{m~{}\!\!n~{}\!\!l}=\nu^{(0)}_{n~{}\!\!l~{}\!\!m}=-\nu^{(0)}_{m~{}\!\!l~{}\!\!n}=-\nu^{(0)}_{l~{}\!\!n~{}\!\!m}=-\nu^{(0)}_{n~{}\!\!m~{}\!\!l}$.
For later convenience, we adopt a manifestly cyclic form such as
$\nu_{l~{}\!\!m~{}\!\!n}$ $=\beta\left\{(E_{1})_{l~{}\!\!n}(E_{2})_{m~{}\!\!n}-(E_{2})_{l~{}\!\!n}(E_{1})_{m~{}\!\!n}\right.+(E_{1})_{m~{}\!\!l}(E_{2})_{n~{}\!\!l}-(E_{2})_{m~{}\!\!l}(E_{1})_{n~{}\!\!l}$
$~{}~{}~{}~{}\left.+(E_{1})_{n~{}\!\!m}(E_{2})_{l~{}\!\!m}-(E_{2})_{n~{}\!\!m}(E_{1})_{l~{}\!\!m}\right\},$ (68)
as a frequency condition. We note that $\nu_{l~{}\!\!m~{}\!\!n}$ is also totally anti-symmetric and satisfies a generalization of Ritz’s combination rule:
$\nu_{l~{}\!\!m~{}\!\!n}=\nu_{l~{}\!\!m~{}\!\!k}+\nu_{l~{}\!\!k~{}\!\!n}+\nu_{k~{}\!\!m~{}\!\!n}.$ (69)
Next, let us derive an equation of motion for variables with three indices based on the condition (68). We consider a periodic motion with the period $T$. Since $\nu_{l~{}\!\!m~{}\!\!n}$ has three indices, we suppose that a dynamical variable containing it as a frequency also owns three indices. Then, the quantities $(A_{1})_{l~{}\!\!m~{}\!\!n}(t)$, $(A_{2})_{l~{}\!\!m~{}\!\!n}(t)$ and $(A_{3})_{l~{}\!\!m~{}\!\!n}(t)$ have a feature that $(A_{1})_{l~{}\!\!m~{}\!\!n}(t+T)=(A_{1})_{l~{}\!\!m~{}\!\!n}(t)$, $(A_{2})_{l~{}\!\!m~{}\!\!n}(t+T)=(A_{2})_{l~{}\!\!m~{}\!\!n}(t)$ , and $(A_{3})_{l~{}\!\!m~{}\!\!n}(t+T)=(A_{3})_{l~{}\!\!m~{}\!\!n}(t)$ and can be described as
$(A_{1})_{l~{}\!\!m~{}\!\!n}(t)=(A_{1})_{l~{}\!\!m~{}\!\!n}(0)e^{2\pi i\nu_{l~{}\!\!m~{}\!\!n}t},~{}~{}(A_{2})_{l~{}\!\!m~{}\!\!n}(t)=(A_{2})_{l~{}\!\!m~{}\!\!n}(0)e^{2\pi i\nu_{l~{}\!\!m~{}\!\!n}t},~{}~{}$ (70)
$(A_{3})_{l~{}\!\!m~{}\!\!n}(t)=(A_{3})_{l~{}\!\!m~{}\!\!n}(0)e^{2\pi i\nu_{l~{}\!\!m~{}\!\!n}t},$
where $\nu_{l~{}\!\!m~{}\!\!n}$ follows the condition (68) and is also expressed by $\nu_{l~{}\!\!m~{}\!\!n}=n_{l~{}\!\!m~{}\!\!n}/T$ ($n_{l~{}\!\!m~{}\!\!n}$ : an integer).
If we require that the product $(A_{1}A_{2}A_{3})_{l~{}\!\!m~{}\!\!n}(t)$ should take the same form as eq. (70), i.e., $(A_{1}A_{2}A_{3})_{l~{}\!\!m~{}\!\!n}(t)=(A_{1}A_{2}A_{3})_{l~{}\!\!m~{}\!\!n}(0)e^{2\pi i\nu_{l~{}\!\!m~{}\!\!n}t}$, $(A_{1})_{l~{}\!\!m~{}\!\!n}(t)$, $(A_{2})_{l~{}\!\!m~{}\!\!n}(t)$ and $(A_{3})_{l~{}\!\!m~{}\!\!n}(t)$ become components of extended matrices whose product is defined by
$(A_{1}A_{2}A_{3})_{l~{}\!\!m~{}\!\!n}(t)\equiv\sum_{k}(A_{1})_{l~{}\!\!m~{}\!\!k}(t)(A_{2})_{l~{}\!\!k~{}\!\!n}(t)(A_{3})_{k~{}\!\!m~{}\!\!n}(t),$ (71)
where we use the generalized rule (69). We refer to three-index variables, which follow the product rule (71), as cubic matrices333 Three-index objects have been introduced to construct a quantum version of the Nambu bracket in Ref. [21]. Their definition of the 3-fold product is different from ours..
Because the components with $l=m$, $m=n$ and/or $n=l$ on $(A_{1})_{l~{}\!\!m~{}\!\!n}(t)$, $(A_{2})_{l~{}\!\!m~{}\!\!n}(t)$ and $(A_{3})_{l~{}\!\!m~{}\!\!n}(t)$ do not depend on time as seen from eqs. (68) and (70), any conserved quantity satisfying $dC_{l~{}\!\!m~{}\!\!n}/dt=0$ can be written as
$C_{l~{}\!\!m~{}\!\!n}=c_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+c_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+c_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m},$ (72)
where $c_{l~{}\!\!m}$, $c_{m~{}\!\!n}$ and $c_{n~{}\!\!l}$ are constants. We refer to the form of cubic matrices (72) as a normal form or a normal cubic matrix [^4] Although a normal matrix is defined by $C_{l~{}\!\!m~{}\!\!n}=\delta_{l~{}\!\!m}c_{m~{}\!\!n}+\delta_{m~{}\!\!n}c_{n~{}\!\!l}+\delta_{n~{}\!\!l}c_{l~{}\!\!m}$ in Refs. [12, 13], we change the definition in order to match a definition of the normal form on generalized matrices with $n(\geq 3)$ indices presented in subsection 4.1. In eq. (72), we add extra factors such as $(1-\delta_{l~{}\!\!n})$, $(1-\delta_{m~{}\!\!l})$ and $(1-\delta_{n~{}\!\!m})$ to meet the condition $C_{l~{}\!\!l~{}\!\!l}=0$. We note that the condition $C_{l~{}\!\!l~{}\!\!l}=0$ is fulfilled without extra factors if $c_{l~{}\!\!m}$, $c_{m~{}\!\!n}$ and $c_{n~{}\!\!l}$ are anti-symmetric..
In a closed system, the Hamiltonians conserve, and hence $(H_{1})_{l~{}\!\!m~{}\!\!n}$ and $(H_{2})_{l~{}\!\!m~{}\!\!n}$ take the normal form:
$(H_{1})_{l~{}\!\!m~{}\!\!n}=(h_{1})_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(h_{1})_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(h_{1})_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m},$ (73)
$(H_{2})_{l~{}\!\!m~{}\!\!n}=(h_{2})_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(h_{2})_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(h_{2})_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m},$ (74)
where $(h_{1})_{l~{}\!\!m}$, $(h_{1})_{m~{}\!\!n}$, $(h_{1})_{n~{}\!\!l}$, $(h_{2})_{l~{}\!\!m}$, $(h_{2})_{m~{}\!\!n}$ and $(h_{2})_{n~{}\!\!l}$ are constants. Using eqs. (68), (69), (70), (71), (73) and (74), we derive the equation:
$\frac{d}{dt}A_{l~{}\!\!m~{}\!\!n}(t)$ $=2\pi i\nu_{l~{}\!\!m~{}\!\!n}A_{l~{}\!\!m~{}\!\!n}(t)$
$=2\pi i\beta\left\{(E_{1})_{l~{}\!\!n}(E_{2})_{m~{}\!\!n}-(E_{2})_{l~{}\!\!n}(E_{1})_{m~{}\!\!n}+(E_{1})_{m~{}\!\!l}(E_{2})_{n~{}\!\!l}-(E_{2})_{m~{}\!\!l}(E_{1})_{n~{}\!\!l}\right.$
$~{}~{}~{}~{}\left.+(E_{1})_{n~{}\!\!m}(E_{2})_{l~{}\!\!m}-(E_{2})_{n~{}\!\!m}(E_{1})_{l~{}\!\!m}\right\}A_{l~{}\!\!m~{}\!\!n}(t)$
$=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2}\right]_{l~{}\!\!m~{}\!\!n}(t),$ (75)
where we choose $\beta=-1/h$, $(E_{1})_{l~{}\!n}=(h_{1})_{l~{}\!n}$, $(E_{2})_{m~{}\!n}=(h_{2})_{m~{}\!n}$ and so on. In eq. (75), $\left[A(t),H_{1},H_{2}\right]_{l~{}\!\!m~{}\!\!n}$ is a triple commutator defined by
$\left[A(t),H_{1},H_{2}\right]_{l~{}\!\!m~{}\!\!n}$ $\equiv\sum_{k}\left\{A(t)_{l~{}\!\!m~{}\!\!k}(H_{1})_{l~{}\!\!k~{}\!\!n}(H_{2})_{k~{}\!\!m~{}\!\!n}+(H_{1})_{l~{}\!\!m~{}\!\!k}(H_{2})_{l~{}\!\!k~{}\!\!n}A(t)_{k~{}\!\!m~{}\!\!n}\right.$
$~{}~{}~{}~{}+(H_{2})_{l~{}\!\!m~{}\!\!k}A(t)_{l~{}\!\!k~{}\!\!n}(H_{1})_{k~{}\!\!m~{}\!\!n}-(H_{2})_{l~{}\!\!m~{}\!\!k}(H_{1})_{l~{}\!\!k~{}\!\!n}A(t)_{k~{}\!\!m~{}\!\!n}$
$~{}~{}~{}~{}\left.-(H_{1})_{l~{}\!\!m~{}\!\!k}A(t)_{l~{}\!\!k~{}\!\!n}(H_{2})_{k~{}\!\!m~{}\!\!n}-A(t)_{l~{}\!\!m~{}\!\!k}(H_{2})_{l~{}\!\!k~{}\!\!n}(H_{1})_{k~{}\!\!m~{}\!\!n}\right\}.$ (76)
In this way, we have arrived at an extension of Heisenberg’s equation of motion:
$\frac{d}{dt}A_{l~{}\!\!m~{}\!\!n}(t)=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2}\right]_{l~{}\!\!m~{}\!\!n},$ (77)
as a counterpart of the Nambu equation (46). It is reconfirmed that the components with $l=m$, $m=n$ and/or $n=l$ on $A_{l~{}\!\!m~{}\!\!n}(t)$ are not independent of time, using eq. (77) and the relation $[B,C_{1},C_{2}]_{l~{}\!\!m~{}\!\!n}=0$ that holds concerning a cubic matrix $B$ with $B_{l~{}\!\!m~{}\!\!n}=0$ for $l\neq m$, $m\neq n$ and $n\neq l$ and arbitrary normal cubic matrices $C_{1}$ and $C_{2}$.
Under the infinitesimal transformation defined by
$\delta_{G}A_{l~{}\!\!m~{}\!\!n}=\varepsilon[A,G_{1},G_{2}]_{l~{}\!\!m~{}\!\!n}=\varepsilon\widetilde{(G_{1}G_{2})}_{l~{}\!\!m~{}\!\!n}A_{l~{}\!\!m~{}\!\!n}$ (78)
with an infinitesimal constant $\varepsilon$ and normal cubic matrices $(G_{1})_{l~{}\!\!m~{}\!\!n}$ and $(G_{2})_{l~{}\!\!m~{}\!\!n}$, using a cubic matrix version of the fundamental identity555 When normal forms play the role of generators, the validity of the fundamental identity is confirmed for generalized matrices with $n(\geq 3)$ indices in subsection 4.2. In eq. (79), the Hamiltonians $H_{1}$ and $H_{2}$ act as generators.:
$[[A,G_{1},G_{2}],H_{1},H_{2}]_{l~{}\!\!m~{}\!\!n}$ (79)
$~{}~{}~{}=[[A,H_{1},H_{2}],G_{1},G_{2}]_{l~{}\!\!m~{}\!\!n}+[A,[G_{1},H_{1},H_{2}],G_{2}]_{l~{}\!\!m~{}\!\!n}+[A,G_{1},[G_{2},H_{1},H_{2}]]_{l~{}\!\!m~{}\!\!n},$
it is shown that the equation is invariant such that
$\frac{d}{dt}A^{\prime}_{l~{}\!\!m~{}\!\!n}=[A^{\prime},H_{1},H_{2}]_{l~{}\!\!m~{}\!\!n},$ (80)
where $A^{\prime}_{l~{}\!\!m~{}\!\!n}=A_{l~{}\!\!m~{}\!\!n}+\delta_{G}A_{l~{}\!\!m~{}\!\!n}$. In eq. (78), $\widetilde{(G_{1}G_{2})}_{l~{}\!\!m~{}\!\!n}$ is defined by
$\widetilde{(G_{1}G_{2})}_{l~{}\!\!m~{}\!\!n}$ $\equiv(g_{1})_{l~{}\!\!n}(g_{2})_{m~{}\!\!n}-(g_{2})_{l~{}\!\!n}(g_{1})_{m~{}\!\!n}+(g_{1})_{m~{}\!\!l}(g_{2})_{n~{}\!\!l}-(g_{2})_{m~{}\!\!l}(g_{1})_{n~{}\!\!l}$
$~{}~{}~{}~{}+(g_{1})_{n~{}\!\!m}(g_{2})_{l~{}\!\!m}-(g_{2})_{n~{}\!\!m}(g_{1})_{l~{}\!\!m},$ (81)
using normal cubic matrices $(G_{1})_{l~{}\!\!m~{}\!\!n}$ and $(G_{2})_{l~{}\!\!m~{}\!\!n}$ given by
$(G_{1})_{l~{}\!\!m~{}\!\!n}=(g_{1})_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(g_{1})_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(g_{1})_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m},$ (82)
$(G_{2})_{l~{}\!\!m~{}\!\!n}=(g_{2})_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(g_{2})_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(g_{2})_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m}.$ (83)
From eqs. (77) and (78), we find that the time evolution of variables is generated by the transformation whose generators are the Hamiltonians.
Let us deal with eq. (77) as the master equation. Then, we find that there is another time-dependent solution such that
$A_{l~{}\!\!m~{}\!\!n}(t)=A_{l~{}\!\!m~{}\!\!n}(0)e^{2\pi i\tilde{\nu}_{l~{}\!\!m~{}\!\!n}t},$ (84)
$\tilde{\nu}_{l~{}\!\!m~{}\!\!n}=\frac{2}{h}\left\{(\tilde{h}_{1})_{n~{}\!\!l}+(\tilde{h}_{1})_{l~{}\!\!m}+(\tilde{h}_{1})_{m~{}\!\!n}\right\},$ (85)
where $(\tilde{h}_{1})_{l~{}\!\!m}$, $(\tilde{h}_{1})_{m~{}\!\!n}$ and $(\tilde{h}_{1})_{n~{}\!\!l}$ are anti-symmetric constants but do not satisfy the combination rule such as eq. (62), i.e., $(\tilde{h}_{1})_{l~{}\!\!m}\neq(\tilde{h}_{1})_{l~{}\!\!n}+(\tilde{h}_{1})_{n~{}\!\!m}$. As seen from eq. (85), $\tilde{\nu}_{l~{}\!\!m~{}\!\!n}$ takes the form of 3-coboundary and satisfies the cocycle condition:
$(\delta\tilde{\nu})_{l~{}\!\!m~{}\!\!n~{}\!\!k}=\tilde{\nu}_{m~{}\!\!n~{}\!\!k}-\tilde{\nu}_{l~{}\!\!n~{}\!\!k}+\tilde{\nu}_{l~{}\!\!m~{}\!\!k}-\tilde{\nu}_{l~{}\!\!m~{}\!\!n}=0.$ (86)
In this solution, $(H_{1})_{l~{}\!\!m~{}\!\!n}$ and $(H_{2})_{l~{}\!\!m~{}\!\!n}$ take the following normal forms:
$(H_{1})_{l~{}\!\!m~{}\!\!n}=(\tilde{h}_{1})_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(\tilde{h}_{1})_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(\tilde{h}_{1})_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m},$ (87)
$(H_{2})_{l~{}\!\!m~{}\!\!n}=(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m}.$ (88)
3.2.1 Fermionic harmonic oscillator
We study a harmonic oscillator relating to fermionic variables as an example of a solution describing eqs. (84) and (85). They are represented by the $3\times 3\times 3$ cubic matrices whose components are given by
$\xi_{l~{}\!\!m~{}\!\!n}(t)=\frac{1}{\sqrt{2}}|\varepsilon_{l~{}\!\!m~{}\!\!n}|e^{-i\omega\varepsilon_{l~{}\!\!m~{}\!\!n}t},~{}~{}\eta_{l~{}\!\!m~{}\!\!n}(t)=-\frac{i}{\sqrt{2}}\varepsilon_{l~{}\!\!m~{}\!\!n}e^{-i\omega\varepsilon_{l~{}\!\!m~{}\!\!n}t},$ (89)
where $l$, $m$ and $n$ run from $1$ to $3$, $\omega$ is the angular frequency, $\varepsilon_{1~{}\!\!2~{}\!\!3}=\varepsilon_{2~{}\!\!3~{}\!\!1}=\varepsilon_{3~{}\!\!1~{}\!\!2}=-\varepsilon_{3~{}\!\!2~{}\!\!1}=-\varepsilon_{2~{}\!\!1~{}\!\!3}=-\varepsilon_{1~{}\!\!3~{}\!\!2}=1$ and others are zero. These variables satisfy the anti-commutation relations:
$\left\{\xi(t),I,\xi(t)\right\}_{l~{}\!\!m~{}\!\!n}=I_{l~{}\!\!m~{}\!\!n},~{}~{}\left\{\eta(t),I,\eta(t)\right\}_{l~{}\!\!m~{}\!\!n}=I_{l~{}\!\!m~{}\!\!n},~{}~{}\left\{\xi(t),I,\eta(t)\right\}_{l~{}\!\!m~{}\!\!n}=0,$ (90)
where $\{A_{1},A_{2},A_{3}\}_{l~{}\!\!m~{}\!\!n}$ is the triple anti-commutator defined by
$\{A_{1},A_{2},A_{3}\}_{l~{}\!\!m~{}\!\!n}$ $\equiv\sum_{k}\left\{(A_{1})_{l~{}\!\!m~{}\!\!k}(A_{2})_{l~{}\!\!k~{}\!\!n}(A_{3})_{k~{}\!\!m~{}\!\!n}+(A_{2})_{l~{}\!\!m~{}\!\!k}(A_{3})_{l~{}\!\!k~{}\!\!n}(A_{1})_{k~{}\!\!m~{}\!\!n}\right.$
$~{}~{}~{}~{}+(A_{3})_{l~{}\!\!m~{}\!\!k}(A_{1})_{l~{}\!\!k~{}\!\!n}(A_{2})_{k~{}\!\!m~{}\!\!n}+(A_{2})_{l~{}\!\!m~{}\!\!k}(A_{1})_{l~{}\!\!k~{}\!\!n}(A_{3})_{k~{}\!\!m~{}\!\!n}$
$~{}~{}~{}~{}\left.+(A_{1})_{l~{}\!\!m~{}\!\!k}(A_{3})_{l~{}\!\!k~{}\!\!n}(A_{2})_{k~{}\!\!m~{}\!\!n}+(A_{3})_{l~{}\!\!m~{}\!\!k}(A_{2})_{l~{}\!\!k~{}\!\!n}(A_{1})_{k~{}\!\!m~{}\!\!n}\right\},$ (91)
and $I$ is a special type of normal form whose components are defined by
$I_{l~{}\!\!m~{}\!\!n}\equiv(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m}.$ (92)
We find that $I$ can play the role of the identity for the components $A_{l~{}\!\!m~{}\!\!n}$ $(l\neq m,m\neq n,n\neq l)$ from the relation $(AII)_{l~{}\!\!m~{}\!\!n}=(IAI)_{l~{}\!\!m~{}\!\!n}=(IIA)_{l~{}\!\!m~{}\!\!n}=A_{l~{}\!\!m~{}\!\!n}$.
Fermionic variables become solutions of the generalized Heisenberg’s equation of motion:
$\frac{d}{dt}\xi_{l~{}\!\!m~{}\!\!n}(t)=\frac{1}{i\hbar}[\xi(t),H_{1},H_{2}]_{l~{}\!\!m~{}\!\!n},~{}~{}\frac{d}{dt}\eta_{l~{}\!\!m~{}\!\!n}(t)=\frac{1}{i\hbar}[\eta(t),H_{1},H_{2}]_{l~{}\!\!m~{}\!\!n},$ (93)
where the Hamiltonians are given by
$(H_{1})_{l~{}\!\!m~{}\!\!n}$ $=i\frac{\hbar\omega}{6}\left[\xi(t),I,\eta(t)\right]_{l~{}\!\!m~{}\!\!n}=\frac{\hbar\omega}{6}\left(\varepsilon_{l~{}\!\!m}\delta_{m~{}\!\!n}+\varepsilon_{m~{}\!\!n}\delta_{n~{}\!\!l}+\varepsilon_{n~{}\!\!l}\delta_{l~{}\!\!m}\right),$ (94)
$(H_{2})_{l~{}\!\!m~{}\!\!n}$ $=I_{l~{}\!\!m~{}\!\!n}$ (95)
with the anti-symmetric quantity $\varepsilon_{l~{}\!\!m}\equiv\sum_{k=1}^{3}\varepsilon_{l~{}\!\!m~{}\!\!k}$. From eqs. (84), (89) and (93), $\tilde{\nu}_{l~{}\!\!m~{}\!\!n}$ is written as
$\tilde{\nu}_{l~{}\!\!m~{}\!\!n}=-\frac{\omega}{2\pi}\varepsilon_{l~{}\!\!m~{}\!\!n}=-\frac{\omega}{6\pi}\left(\varepsilon_{n~{}\!\!l}+\varepsilon_{l~{}\!\!m}+\varepsilon_{m~{}\!\!n}\right),$ (96)
and it satisfies the cocycle condition $(\delta\tilde{\nu})_{l~{}\!\!m~{}\!\!n~{}\!\!k}=0$. Since $\xi_{l~{}\!\!m~{}\!\!n}(t)$ and $\eta_{l~{}\!\!m~{}\!\!n}(t)$ follow the equations:
$\frac{d}{dt}\xi_{l~{}\!\!m~{}\!\!n}(t)=\omega{\eta}_{l~{}\!\!m~{}\!\!n}(t),~{}~{}\frac{d}{dt}\eta_{l~{}\!\!m~{}\!\!n}(t)=-\omega{\xi}_{l~{}\!\!m~{}\!\!n}(t),$ (97)
they obey the equation for simple harmonic motion:
$\frac{d^{2}}{dt^{2}}\xi_{l~{}\!\!m~{}\!\!n}(t)=-\omega^{2}{\xi}_{l~{}\!\!m~{}\!\!n}(t),~{}~{}\frac{d^{2}}{dt^{2}}\eta_{l~{}\!\!m~{}\!\!n}(t)=-\omega^{2}{\eta}_{l~{}\!\!m~{}\!\!n}(t).$ (98)
The same system is described by the cubic matrices interpreted as the ladder operators:
$C_{l~{}\!\!m~{}\!\!n}(t)\equiv\frac{1}{\sqrt{2}}\left(\xi_{l~{}\!\!m~{}\!\!n}(t)+i\eta_{l~{}\!\!m~{}\!\!n}(t)\right)=\frac{1}{2}\left(|\varepsilon_{l~{}\!\!m~{}\!\!n}|+\varepsilon_{l~{}\!\!m~{}\!\!n}\right)e^{-i\omega t},~{}~{}$ (99)
$C^{\dagger}_{l~{}\!\!m~{}\!\!n}(t)\equiv\frac{1}{\sqrt{2}}\left(\xi_{l~{}\!\!m~{}\!\!n}(t)-i\eta_{l~{}\!\!m~{}\!\!n}(t)\right)=\frac{1}{2}\left(|\varepsilon_{l~{}\!\!m~{}\!\!n}|-\varepsilon_{l~{}\!\!m~{}\!\!n}\right)e^{i\omega t}.$ (100)
These variables satisfy the anti-commutation relations:
$\{C(t),I,C^{\dagger}(t)\}_{l~{}\!\!m~{}\!\!n}=I_{l~{}\!\!m~{}\!\!n},~{}~{}\{C(t),I,C(t)\}_{l~{}\!\!m~{}\!\!n}=0,~{}~{}\{C^{\dagger}(t),I,C^{\dagger}(t)\}_{l~{}\!\!m~{}\!\!n}=0,$ (101)
and become solutions of the generalized Heisenberg’s equation of motion:
$\frac{d}{dt}C_{l~{}\!\!m~{}\!\!n}(t)=\frac{1}{i\hbar}[C(t),H_{1},H_{2}]_{l~{}\!\!m~{}\!\!n},~{}~{}\frac{d}{dt}C^{\dagger}_{l~{}\!\!m~{}\!\!n}(t)=\frac{1}{i\hbar}[C^{\dagger}(t),H_{1},H_{2}]_{l~{}\!\!m~{}\!\!n},$ (102)
where the Hamiltonians are written by
$(H_{1})_{l~{}\!\!m~{}\!\!n}=\frac{\hbar\omega}{6}[C^{\dagger}(t),I,C(t)]_{l~{}\!\!m~{}\!\!n},~{}~{}(H_{2})_{l~{}\!\!m~{}\!\!n}=I_{l~{}\!\!m~{}\!\!n}.$ (103)
Since $C_{l~{}\!\!m~{}\!\!n}(t)$ and $C^{\dagger}_{l~{}\!\!m~{}\!\!n}(t)$ follow the equations:
$\frac{d}{dt}C_{l~{}\!\!m~{}\!\!n}(t)=-i\omega C_{l~{}\!\!m~{}\!\!n}(t),~{}~{}\frac{d}{dt}C^{\dagger}_{l~{}\!\!m~{}\!\!n}(t)=i\omega C^{\dagger}_{l~{}\!\!m~{}\!\!n}(t),$ (104)
they also obey the equation for simple harmonic motion.
4 Generalized Heisenberg dynamics of variables with $n$ indices
4.1 Generalized matrix mechanics
Let us start from a generalization of eq. (54):
$\nu=\frac{\partial(E_{1},E_{2},\cdots,E_{n-1})}{\partial(J_{1},J_{2},\cdots,J_{n-1})}=\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\frac{\partial E_{a_{1}}}{\partial J_{1}}\frac{\partial E_{a_{2}}}{\partial J_{2}}\cdots\frac{\partial E_{a_{n-1}}}{\partial J_{n-1}},$ (105)
where sgn(P) is $+1$ and $-1$ for even and odd permutations among the set of subscripts $(a_{1},a_{2},\cdots,a_{n-1})$, respectively, based on $+1$ for $(a_{1},a_{2},\cdots,a_{n-1})=(1,2,\cdots,n-1)$, and each $a_{k}$ $(k=1,2,\cdots,n-1)$ runs from $1$ to $n-1$. In the following, we construct a mechanics of quantities with $n$ indices, using the correspondence principle, or tackle a puzzle like a cryptarithm shown in Figure 3.
Figure 3: Classical theory and its quantum (or quantum analog of) theory.
We assume that there is a one-to-one correspondence between $E_{a_{k}}$ in eq. (105) and some discrete values $(E_{a_{k}})_{l_{r}~{}\!\!l_{r^{\prime}}}$ ($l_{r}$, $l_{r^{\prime}}$ : natural numbers, $r,r^{\prime}=1,2,\cdots,n$), and these discrete values are anti-symmetric as
$(E_{a_{k}})_{l_{r^{\prime}}~{}\!\!l_{r}}=-(E_{a_{k}})_{l_{r}~{}\!\!l_{r^{\prime}}}$ (106)
and follow the combination rule:
$(E_{a_{k}})_{l_{r}~{}\!\!l_{r^{\prime}}}=(E_{a_{k}})_{l_{r}~{}\!\!l_{r^{\prime\prime}}}+(E_{a_{k}})_{l_{r^{\prime\prime}}~{}\!\!l_{r^{\prime}}}=(E_{a_{k}})_{l_{r}~{}\!\!l_{r^{\prime\prime}}}-(E_{a_{k}})_{l_{r^{\prime}}~{}\!\!l_{r^{\prime\prime}}},$ (107)
where $l_{r^{\prime\prime}}$ is also a natural number.
By identifying ${\partial E_{a_{k}}}/{\partial J_{b}}$ with ${(E_{a_{k}})_{l_{b}~{}\!\!l_{n}}}/{\varDelta l_{b}}$ ($l_{n}=l_{b}-\varDelta l_{b}$, $\varDelta l_{b}$ : a natural number, $b=1,2,\cdots,n-1$), we obtain the relation such as
$(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}=\beta\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{n-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{n}},$ (108)
where $\beta$ is a constant, and $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ is regarded as a counterpart of $\nu\varDelta l_{1}\varDelta l_{2}\cdots\varDelta l_{n-1}$. From eq. (108), we see that $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ is anti-symmetric under the exchange of $n-1$ indices $(l_{1},l_{2},\cdots,l_{n-1})$.
Using eqs. (106), (107) and (108), we derive the relation666 The details of derivation of this relation and most other equations are presented in Appendix A.:
$(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=\sum_{i=1}^{n}(-1)^{n-i}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}},$ (109)
where the hatted index $\hat{l}_{i}$ means that $l_{i}$ is omitted, and $l_{n+1}$ is a natural number. Eq. (109) implies that $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ is a $n$-cocycle, i.e., $(\delta\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}=0$. Here, $\delta$ is a coboundary operator defined by
$(\delta\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}=\sum_{i=1}^{n+1}(-1)^{i+1}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n+1}}.$ (110)
Let us require that $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ should have a cyclic property among indices777 The frequency $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ given by eq. (108) also have a cyclic property in a hidden way, as shown in Appendix A, and hence we find that it is totally anti-symmetric under the exchange of $(l_{1},l_{2},\cdots,l_{n})$. such as
$(\nu_{n})_{l_{p+1}~{}\!\!l_{p+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{p}}=(-1)^{(n-1)p}(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (111)
in a manifest way. Here, we use the notation such as $l_{p}=l_{q}$ with $p=q$ (mod $n$) and do it hereafter. The following type of frequency possesses a manifestly cyclic symmetry
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=\sum_{k=1}^{n}(-1)^{(n-1)k}(\nu_{n}^{(0)})_{l_{k+1}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{k}},$ (112)
and then it is also totally anti-symmetric under the exchange of $(l_{1},l_{2},\cdots,l_{n})$.
Using eqs. (108) and (112), we obtain the relation:
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}=\beta\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}.$ (113)
We find that eq. (109) holds for $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ and $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ also becomes $n$-cocycle and satisfies a generalization of Ritz’s combination rule:
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}+(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}+\cdots+(\nu_{n})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}.$ (114)
Next, we derive an equation of motion for quantities with $n$ indices based on the condition (113). We consider a periodic motion with the period $T$. Because $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ has $n$ indices, we suppose that a dynamical variable containing it as a frequency also possesses $n$ indices. Then, the quantities $(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ $(a=1,2,\cdots,n)$ have a feature that $(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t+T)=(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ and can be described as
$(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(0)e^{2\pi i(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}t},$ (115)
where $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ follows the condition (113) and is also written by $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=n_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}/T$ ($n_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ : an integer).
If we require that the product $(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ should take the same form as eq. (115), i.e., $(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(0)e^{2\pi i(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}t}$, $(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ become components of extended matrices whose product is defined by888 This definition of the $n$-fold product is the same as that in Ref. [22].
$(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)\equiv\sum_{l_{n+1}}(A_{1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}(t)(A_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}~{}\!\!l_{n}}(t)\cdots(A_{n})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t),$ (116)
where we use eq. (114). We refer to $n(\geq 3)$-index variables, which follow the product rule (116), as generalized matrices or $n$-th power matrices.
Because the components with $l_{r}=l_{r^{\prime}}$ on $(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ do not depend on time as seen from eqs. (112) and (115), any conserved quantity satisfying $dC_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}/dt=0$ can be described as a generalized matrix whose non-vanishing components are given by
$C_{l_{1}~{}\!\!\cdots~{}\!\!l_{i}~{}\!\!\cdots~{}\!\!l_{j-1}~{}\!\!l_{i}~{}\!\!l_{j+1}~{}\!\!\cdots~{}\!\!l_{n}}=\sum_{u(\neq i,j)}c_{l_{u}~{}\!\!l_{i}},$ (117)
where $c_{l_{u}~{}\!\!l_{i}}$ are constants, $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, $u$ runs from 1 to $n$ except for $i$ and $j$, and the $n-1$ indices $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different. We note that other components are zero. We refer to this type of generalized matrix as a normal form, a normal generalized matrix or an $n$-th power normal matrix.
Then, in a closed system, the Hamiltonians $(H_{a_{k}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ ($k=1,2\cdots,n-1$) conserve and take the normal forms whose non-vanishing components are given by
$(H_{a_{k}})_{l_{1}~{}\!\!\cdots~{}\!\!l_{i}~{}\!\!\cdots~{}\!\!l_{j-1}~{}\!\!l_{i}~{}\!\!l_{j+1}~{}\!\!\cdots~{}\!\!l_{n}}=\sum_{u(\neq i,j)}(h_{a_{k}})_{l_{u}~{}\!\!l_{i}},$ (118)
where $(h_{a_{k}})_{l_{u}~{}\!\!l_{i}}$ are constants, $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, $u$ runs from 1 to $n$ except for $i$ and $j$, and the $n-1$ indices $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different.
Using eqs. (113), (114), (115), (116) and (118), we derive the equation:
$\frac{d}{dt}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ $=2\pi i(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$
$=2\pi i\beta\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}$
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$
$=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (119)
by choosing $\beta$, $(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}$ and $(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ as
$\beta=\frac{\gamma}{h}(-1)^{\frac{(n-2)(n-3)}{2}+1},~{}~{}$ (120)
$(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}=(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}~{}~{}(r=1,\cdots,k-1),~{}~{}$ (121)
$(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}=(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}~{}~{}(r^{\prime}=k,\cdots,n-1),$ (122)
where $\gamma=1$ for an odd number $n(\geq 3)$ and $\gamma=n-2$ for an even number $n(\geq 4)$. Using eqs. (113), (120), (121) and (122), the frequency is represented by
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ $=\frac{\gamma}{h}(-1)^{\frac{(n-2)(n-3)}{2}+1}\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}$
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}.$ (123)
In eq. (119), $\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ is the $n$-fold commutator defined by
$\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ (124)
$~{}~{}~{}~{}\equiv\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\sum_{l_{n+1}}\left\{A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}(H_{a_{1}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}(H_{a_{2}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-3}~{}\!\!l_{n+1}~{}\!\!l_{n-1}~{}\!\!l_{n}}\right.$
$~{}~{}~{}~{}~{}~{}~{}\cdots(H_{a_{n-1}})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}-(H_{a_{1}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}(H_{a_{2}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-3}~{}\!\!l_{n+1}~{}\!\!l_{n-1}~{}\!\!l_{n}}$
$~{}~{}~{}~{}~{}~{}~{}\cdots(H_{a_{n-1}})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}+\cdots+(-1)^{n-1}(H_{a_{1}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}(H_{a_{2}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}$
×$~{}~{}~{}~{}~{}~{}~{}\left.\times(H_{a_{3}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-3}~{}\!\!l_{n+1}~{}\!\!l_{n-1}~{}\!\!l_{n}}\cdots(H_{a_{n-1}})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}A(t)_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\right\}.$
In this way, we have arrived at a generalization of Heisenberg’s equation of motion:
$\frac{d}{dt}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}.$ (125)
In the same way as the dynamics of cubic matrices, it is reconfirmed that the components with identical indices on $A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ are not independent of time, using eq. (125) and the relation $[B,C_{1},\cdots,C_{n-2},C_{n-1}]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ that holds concerning a generalized matrix $B$ whose components with all different indices vanish, i.e., $B_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ ($l_{i}\neq l_{j}$ for all $i$ and $j$) and arbitrary normal cubic matrices $C_{a}$ $(a=1,2,\cdots,n-1)$. The validity of $[B,C_{1},\cdots,C_{n-2},C_{n-1}]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ is understood from the following reason. For $B$ and $C_{a}$, we obtain the relation:
$(BC_{1}\cdots C_{n-2}C_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{l_{n+1}}B_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}(C_{1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}\cdots$
×$~{}~{}~{}~{}~{}\times(C_{n-2})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(C_{n-1})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ (126)
from the fact that there exists no instance in which each of indices $(l_{1},l_{2},\cdots,l_{n-2},l_{n+1},l_{n})$, $\cdots$, $(l_{1},l_{n+1},l_{3},\cdots,l_{n})$ and $(l_{n+1},l_{2},\cdots,l_{n})$ contains exactly two identical indices, other indices are all different, $(l_{1},l_{2},\cdots,l_{n-1},l_{n+1})$ contains two or more than two identical indices, and furthermore, the identical index is shared between them.
Here, we give comments on the equation (125).
One is a symmetry in the equation. Eq. (125) is invariant under the transformation of Hamiltonians $(H_{a_{k}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\to(H^{\prime}_{a_{k}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ whose values given by eq. (118) are transformed as
$(h_{a_{k}})_{l_{u}~{}\!\!l_{i}}\to(h^{\prime}_{a_{k}})_{l_{u}~{}\!\!l_{i}}=(h_{a_{k}})_{l_{u}~{}\!\!l_{i}}+c_{a_{k}},$ (127)
where $c_{a_{k}}$ are constants. Actually, it is shown that the following relation holds on,
$\left[A(t),H^{\prime}_{1},H^{\prime}_{2},\cdots,H^{\prime}_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ (128)
using the relations:
$\left[A(t),H_{a_{1}},H_{a_{2}},\cdots,H_{a_{n-2}},I\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=-h(\delta\nu^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$
$=-h\sum_{i=1}^{n}(-1)^{i+1}(\nu^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots l_{n}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=0,$ (129)
$\left[A(t),H_{a_{1}},\cdots,H_{a_{k}},I,\cdots,I\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=0~{}~{}~{}~{}(k=1,\cdots,n-3),$ (130)
$\left[A(t),I,\cdots,I\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=0,$ (131)
where $I$ is a specific type of normal generalized matrix whose non-vanishing components are given by
$I_{l_{1}~{}\!\!\cdots~{}\!\!l_{i}~{}\!\!\cdots~{}\!\!l_{j-1}~{}\!\!l_{i}~{}\!\!l_{j+1}~{}\!\!\cdots~{}\!\!l_{n}}=1,$ (132)
and eqs.(130) and (131) are understood from the definition of the $n$-fold commutator. In eq. (132), $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, and $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different. In eq. (129), $(\nu^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n-1}}$ is given by
$(\nu^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n-1}}$ $=\frac{\gamma}{h}(-1)^{\frac{(n-3)(n-4)}{2}+1}\sum_{k=1}^{n-1}(-1)^{n-1-k}\sum_{(a_{1},a_{2},\cdots,a_{n-2})}{\rm sgn(P)}$
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-2}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (133)
with $\gamma=1$ for an odd number $n(\geq 3)$ and $\gamma=n-2$ for an even number $n(\geq 4)$. In eq. (129), the combination rule $(h_{a_{k}})_{l_{s}~{}\!\!l_{t}}=(h_{a_{k}})_{l_{s}~{}\!\!l_{u}}+(h_{a_{k}})_{l_{u}~{}\!\!l_{t}}=(h_{a_{k}})_{l_{s}~{}\!\!l_{u}}-(h_{a_{k}})_{l_{t}~{}\!\!l_{u}}$ is used to show $(\delta\nu^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$. We notice that $I$ behaves like the identity, because the following relation holds
$(AI\cdots I)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=(IAI\cdots I)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=\cdots=(I\cdots IA)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ (134)
for the components $A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ with all different indices.
The other is a solution of eq. (125), in the case that eq. (125) is regarded as a master equation. There is another solution such that
$A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(0)e^{2\pi i(\tilde{\nu}_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}t},$ (135)
$(\tilde{\nu}_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}=(\delta\tilde{\nu}^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}=\sum_{i=1}^{n}(-1)^{i+1}(\tilde{\nu}^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots l_{n}},$ (136)
where $(\tilde{\nu}^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n-1}}$ is given by
$(\tilde{\nu}^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n-1}}$ $=\frac{{\gamma}}{h}(-1)^{\frac{(n-3)(n-4)}{2}+1}\sum_{k=1}^{n-1}(-1)^{n-1-k}\sum_{(a_{1},a_{2},\cdots,a_{n-2})}{\rm sgn(P)}$
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(\tilde{h}_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-2}(\tilde{h}_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (137)
with ${\gamma}=1$ for an odd number $n(\geq 3)$ and ${\gamma}=n-2$ for an even number $n(\geq 4)$. In eq. (137), $(\tilde{h}_{a_{r}})_{l_{s}~{}\!\!l_{t}}$ are anti-symmetric constants but do not satisfy the combination rule, i.e., $(\tilde{h}_{a_{k}})_{l_{s}~{}\!\!l_{t}}\neq(\tilde{h}_{a_{k}})_{l_{s}~{}\!\!l_{u}}+(\tilde{h}_{a_{k}})_{l_{u}~{}\!\!l_{t}}$. In this case, Hamiltonians are normal forms whose non-vanishing components are given by
$(H_{a_{k}})_{l_{1}~{}\!\!\cdots~{}\!\!l_{i}~{}\!\!\cdots~{}\!\!l_{j-1}~{}\!\!l_{i}~{}\!\!l_{j+1}~{}\!\!\cdots~{}\!\!l_{n}}=\sum_{u(\neq i,j)}(\tilde{h}_{a_{k}})_{l_{u}~{}\!\!l_{i}}~{}~{}~{}~{}(k=1,2,\cdots,n-2),$ (138)
$(H_{n-1})_{l_{1}~{}\!\!\cdots~{}\!\!l_{i}~{}\!\!\cdots~{}\!\!l_{j-1}~{}\!\!l_{i}~{}\!\!l_{j+1}~{}\!\!\cdots~{}\!\!l_{n}}=1.$ (139)
From eq. (136), we find that $(\tilde{\nu}_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ is $n$-coboundary and satisfies the cocycle condition as seen from the nilpotency of coboundary operator, i.e.,
$(\delta\tilde{\nu}_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n+1}}=(\delta^{2}\tilde{\nu}^{\prime}_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n+1}}=0.$ (140)
4.2 Discretization of Jacobian
Eq. (125) resembles the Nambu equation for $A=A(x_{1},x_{2},\cdots,x_{n})$:
$\frac{dA}{dt}=\{A,H_{1},H_{2},\cdots,H_{n-1}\}_{\rm NB},$ (141)
where $x_{i}$ $(i=1,2,\cdots,n)$ are canonical $n$-plets, $\{A,H_{1},H_{2},\cdots,H_{n-1}\}_{\rm NB}$ is the Nambu bracket and $\{A_{1},A_{2},A_{3},\cdots,A_{n}\}_{\rm NB}$ is defined by
$\{A_{1},A_{2},A_{3},\cdots,A_{n}\}_{\rm NB}$ $\equiv\frac{\partial(A_{1},A_{2},A_{3},\cdots,A_{n})}{\partial(x_{1},x_{2},x_{3},\cdots,x_{n})}$
$=\sum_{(a_{1},a_{2},\cdots,a_{n})}{\rm sgn(P)}\frac{\partial A_{a_{1}}}{\partial x_{1}}\frac{\partial A_{a_{2}}}{\partial x_{2}}\frac{\partial A_{a_{3}}}{\partial x_{3}}\cdots\frac{\partial A_{a_{n}}}{\partial x_{n}},$ (142)
where sgn(P) is $+1$ and $-1$ for even and odd permutations among the set of subscripts $(a_{1},a_{2},\cdots,a_{n})$, respectively, based on $+1$ for $(a_{1},a_{2},\cdots,a_{n})=(1,2,\cdots,n)$, and each $a_{k}$ $(k=1,2,\cdots,n)$ runs from $1$ to $n$. Hence, we conjecture that the generalized Heisenberg’s equation of motion (125) can be a counterpart of the Nambu equation and the multiple commutator defined by
$\left[A_{1},A_{2},\cdots,A_{n}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ (143)
$~{}~{}~{}~{}\equiv\sum_{(a_{1},a_{2},\cdots,a_{n})}{\rm sgn(P)}\sum_{l_{n+1}}(A_{a_{1}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}(A_{a_{2}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}\cdots(A_{a_{n}})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
can become a candidate of discrete or quantum version of the Nambu bracket or the Jacobian.
Let us study whether the Nambu bracket and the multiple commutator possess common features or not. The basic features of the Nambu bracket are listed as follows,
$\{A_{a_{1}},A_{a_{2}},\cdots,A_{a_{n}}\}_{\rm NB}={\rm sgn(P)}\{A_{1},A_{2},\cdots,A_{n}\}_{\rm NB}~{}~{}~{}~{}~{}({\rm skew~{}symmetry}),$ (144)
$\{A_{1}+B_{1},A_{2},\cdots,A_{n}\}_{\rm NB}=\{A_{1},A_{2},\cdots,A_{n}\}_{\rm NB}+\{B_{1},A_{2},\cdots,A_{n}\}_{\rm NB}~{}~{}~{}~{}~{}({\rm linearity}),$ (145)
$\{\{A_{1},A_{2},\cdots,A_{n}\}_{\rm NB},B_{1},B_{2},\cdots,B_{n-1}\}_{\rm NB}$
$~{}~{}~{}~{}~{}=\sum_{i=1}^{n}\{A_{1},A_{2},\cdots,\{A_{i},B_{1},B_{2},\cdots,B_{n-1}\}_{\rm NB},\cdots,A_{n}\}_{\rm NB}~{}~{}~{}~{}~{}({\rm fundamental~{}identity}),$ (146)
$\{A_{1}A_{2}\cdots A_{n},B_{1},B_{2},\cdots,B_{n-1}\}_{\rm NB}=\sum_{i=1}^{n}A_{1}A_{2}\cdots\{A_{i},B_{1},B_{2},\cdots,B_{n-1}\}_{\rm NB}\cdots A_{n}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}({\rm derivation~{}rule}).$ (147)
The multiple commutator of generalized matrices has the skew symmetry and the linearity, but does not necessarily satisfy the fundamental identity and the derivation rule. When $B_{j}$ $(j=1,2,\cdots,n-1)$ are normal generalized matrices and the eigenvalues $(f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ satisfy the cocycle condition $(\delta f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}=0$, we find that the multiple commutator of generalized matrices follows the fundamental identity and the derivation rule, in the following way. Here, the eigenvalues $(f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ are specified by the eigenvalue equation:
$\left[A,B_{1},B_{2},\cdots,B_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=(f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ (148)
and the cocycle condition is denoted by
$(\delta f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}=(-1)^{n+1}\left\{(f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}-\sum_{i=1}^{n}(-1)^{n-i}(f_{n})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}\right\}=0.$ (149)
Using eq. (148), the left-hand side and the right-hand side of the fundamental identity can be calculated as
$\left[[A_{1},A_{2},\cdots,A_{n}],B_{1},B_{2},\cdots,B_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=(f_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}[A_{1},A_{2},\cdots,A_{n}]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (150)
$\sum_{i=1}^{n}\left[A_{1},\cdots,[A_{i},B_{1},B_{2},\cdots,B_{n-1}],\cdots,A_{n}\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\sum_{i=1}^{n}(f_{n})_{l_{1}~{}\!\!\cdots~{}\!\!l_{i-1}~{}\!\!l_{n+1}~{}\!\!l_{i+1}~{}\!\!\cdots~{}\!\!l_{n}}\left[A_{1},A_{2},\cdots,A_{n}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\sum_{i=1}^{n}(-1)^{n-i}(f_{n})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}\left[A_{1},A_{2},\cdots,A_{n}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (151)
respectively, and these agree with if the condition (149) is met. The same holds on for the derivation rule.
5 Conclusions and discussions
Inspired by the fact that Heisenberg’s matrix mechanics was derived from Hamiltonian mechanics under the guidance of the correspondence principle, we have investigated a mechanics involving discrete variables, starting from Nambu mechanics. As a result, we reconstructed a new framework of mechanics for dynamical variables expressed in terms of generalized matrices, which can be regarded as an extension of matrix mechanics. The fundamental equation of this mechanics corresponds to a generalized version of Heisenberg’s equation of motion, formulated using an $n$-fold commutator. This equation can also be interpreted as the counterpart to the Nambu equation associated with canonical $n$-plets. Based on this interpretation, we reconfirmed that the multiple commutator involving generalized matrices can serve as a quantum or discrete version of the Nambu bracket, namely the Jacobian.
In what follows, we discuss the generalized matrix mechanics and possible applications of the generalized matrices.
In quantum mechanics, the matrix elements in matrix mechanics are interpreted as probability amplitudes representing transitions between quantum states. At present, it is unclear whether the generalized matrix elements with $n$ indices admit a similar interpretation, or whether the extended version of matrix mechanics can be applied to real physical systems. This suggests the need for further investigation into the understanding and application of generalized matrices, beyond the framework of a mere extension of Heisenberg’s matrix mechanics or a counterpart to Nambu mechanics.
Next, regarding the use of generalized matrices and the discrete version of the Nambu bracket expressed via the multiple commutator: in quantum mechanics, spin variables are represented using matrices. Therefore, if an extended version of spin exists, it is expected to be representable using generalized matrices. In fact, in Ref. [25], a generalization of spin algebra using generalized matrices has been proposed. Furthermore, in that work, uncertainty relations for spacetime have been derived by using the triple commutator involving spacetime coordinates expressed via cubic matrices. It is also conceivable that uncertainty relations for higher-dimensional spacetime can be derived using the $n$-fold commutator.
Finally, generalized matrices may prove useful in formulating the quantum theory of extended objects. The action integral in string theory includes the area of the worldsheet, and this area is expressed using a two-dimensional Jacobian [26, 27]. Similarly, $p$-dimensional extended objects are described using $(p+1)$-dimensional Jacobians. By discretizing the Jacobian, it may become possible to microscopically formulate various branes including membrane [28], which could, in turn, be applicable to the formulation of M-theory999 Using Lie 3-algebra [29], an effective theory for multiple M2 branes called the BLG model has been constructed [30, 31, 32, 33]. See Ref. [34] for extensive reviews of $n$-nary algebra and its applications.. Recently, there have been attempts relating to many-index objects, for instance, to regularize the Nambu brackets using $2n$-index objects [35] and to construct gauge theories using a new type of non-associative algebra that includes both ordinary matrices and cubic matrices [36]. These developments suggest that there may still be room for various extensions and applications.
Acknowledgments
This work was supported in part by scientific grants from the Ministry of Education, Culture, Sports, Science and Technology under Grant No. 22K03632.
Appendix A Details of derivation
In this appendix, details of derivation of eq. (109), $(\delta\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}=0$ and so on are provided, and it is shown that variables with $n$ totally anti-symmetric indices satisfy the generalized Heisenberg equation (125).
Starting from eq. (108), eq. (109) is derived as follows,
$(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{n-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{n}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{1}~{}\!\!l_{n}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n}}\cdots(E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}((E_{a_{1}})_{l_{1}~{}\!\!l_{n+1}}-(E_{a_{1}})_{l_{n}~{}\!\!l_{n+1}})((E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}-(E_{a_{2}})_{l_{n}~{}\!\!l_{n+1}})$
××$~{}~{}~{}~{}\times\cdots\times((E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n+1}}-(E_{a_{n-1}})_{l_{n}~{}\!\!l_{n+1}})$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\left\{(E_{a_{1}})_{l_{1}~{}\!\!l_{n+1}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}\cdots(E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n+1}}\right.$
$~{}~{}~{}~{}-(E_{a_{1}})_{l_{1}~{}\!\!l_{n+1}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}\cdots(E_{a_{n-2}})_{l_{n-2}~{}\!\!l_{n+1}}(E_{a_{n-1}})_{l_{n}~{}\!\!l_{n+1}}$
$~{}~{}~{}~{}-(E_{a_{1}})_{l_{1}~{}\!\!l_{n+1}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}\cdots(E_{a_{n-2}})_{l_{n}~{}\!\!l_{n+1}}(E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n+1}}$
$~{}~{}~{}~{}\left.-\cdots-(E_{a_{1}})_{l_{n}~{}\!\!l_{n+1}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}\cdots(E_{a_{n-2}})_{l_{n-2}~{}\!\!l_{n+1}}(E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n+1}}\right\}$
$=(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}-\sum_{i=1}^{n-1}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots l_{i-1}~{}\!\!l_{n}~{}\!\!l_{i+1}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}$
$=(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}-\sum_{i=1}^{n-1}(-1)^{n-1-i}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}$
$=\sum_{i=1}^{n}(-1)^{n-i}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}},$ (152)
where we use eqs. (106) and (107) and the feature that terms including both $(E_{a_{r}})_{l_{n}~{}\!\!l_{n+1}}$ and $(E_{a_{r^{\prime}}})_{l_{n}~{}\!\!l_{n+1}}$ ($a_{r}\neq a_{r^{\prime}}$) disappear because they are canceled out due to totally antisymmetric property regarding $(a_{1},a_{2},\cdots,a_{n-1})$.
For reference, the frequencies $(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots l_{i-1}~{}\!\!l_{n}~{}\!\!l_{i+1}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}$ and $(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}$ are expressed as
$(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots l_{i-1}~{}\!\!l_{n}~{}\!\!l_{i+1}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{1}~{}\!\!l_{n+1}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}\cdots$
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times(E_{a_{i-1}})_{l_{i-1}~{}\!\!l_{n+1}}(E_{a_{i}})_{l_{n}~{}\!\!l_{n+1}}(E_{a_{i+1}})_{l_{i+1}~{}\!\!l_{n+1}}\cdots(E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n+1}},$ (153)
$(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{1}~{}\!\!l_{n+1}}(E_{a_{2}})_{l_{2}~{}\!\!l_{n+1}}\cdots$
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times(E_{a_{i-1}})_{l_{i-1}~{}\!\!l_{n+1}}(E_{a_{i}})_{l_{i+1}~{}\!\!l_{n+1}}\cdots(E_{a_{n-2}})_{l_{n-1}~{}\!\!l_{n+1}}(E_{a_{n-1}})_{l_{n}~{}\!\!l_{n+1}},$ (154)
respectively.
It is shown that $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ satisfies the cocycle condition $(\delta\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}=0$ as follows,
$(\delta\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}$ $=\sum_{i=1}^{n+1}(-1)^{i+1}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n+1}}=(-1)^{n}\sum_{i=1}^{n+1}(-1)^{n-i-1}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n+1}}$
$=(-1)^{n}\left\{-\sum_{i=1}^{n}(-1)^{n-i}(\nu_{n}^{(0)})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}+(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\right\}=0,$ (155)
where we use eq. (109).
It is shown that the frequency (112) have a cyclic property (111) as follows,
$(\nu_{n})_{l_{p+1}~{}\!\!l_{p+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{p}}$ $=\sum_{k=1}^{n}(-1)^{(n-1)k}(\nu_{n}^{(0)})_{l_{p+k+1}~{}\!\!l_{p+k+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{p+k}}$
$=(-1)^{(n-1)p}\sum_{k=1}^{n}(-1)^{(n-1)(p+k)}(\nu_{n}^{(0)})_{l_{p+k+1}~{}\!\!l_{p+k+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{p+k}}$
$=(-1)^{(n-1)p}\sum_{k=1}^{n}(-1)^{(n-1)k}(\nu_{n}^{(0)})_{l_{k+1}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{k}}$
$=(-1)^{(n-1)p}(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}.$ (156)
Next, let us derive eq. (113). Using eq. (108), we obtain the relation:
$(\nu_{n}^{(0)})_{l_{k+1}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{k}}$ $={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{n-1}(E_{a_{r}})_{l_{r+k}~{}\!\!l_{k}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{k+1}~{}\!\!l_{k}}(E_{a_{2}})_{l_{k+2}~{}\!\!l_{k}}\cdots(E_{a_{n-k}})_{l_{n}~{}\!\!l_{k}}$
×$~{}~{}~{}~{}\times(E_{a_{n-k+1}})_{l_{1}~{}\!\!l_{k}}(E_{a_{n-k+2}})_{l_{2}~{}\!\!l_{k}}\cdots(E_{a_{n-1}})_{l_{k-1}~{}\!\!l_{k}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(-1)^{(n-k)(k-1)}(E_{a_{1}})_{l_{1}~{}\!\!l_{k}}(E_{a_{2}})_{l_{2}~{}\!\!l_{k}}\cdots$
×$~{}~{}~{}~{}\times(E_{a_{k-1}})_{l_{k-1}~{}\!\!l_{k}}(E_{a_{k}})_{l_{k+1}~{}\!\!l_{k}}(E_{a_{k+1}})_{l_{k+2}~{}\!\!l_{k}}\cdots(E_{a_{n-1}})_{l_{n}~{}\!\!l_{k}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(-1)^{(n-k)(k-1)}$
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},$ (157)
where the factor $(-1)^{(n-k)(k-1)}$ appears from the $(n-k)(k-1)$ exchange of the indices on $E_{a_{r}}$ from $(a_{1},a_{2},\cdots,a_{n-k},a_{n-k+1},a_{n-k+2},\cdots,a_{n-1})$ to $(a_{k},a_{k+1},\cdots,a_{n-1},a_{1},a_{2},\cdots,a_{k-1})$. By inserting eq. (157) into eq. (112), eq. (113) is obtained as follows,
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ $\equiv\sum_{k=1}^{n}(-1)^{(n-1)k}(\nu_{n}^{(0)})_{l_{k+1}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{k}}$
$=\sum_{k=1}^{n}(-1)^{(n-1)k}{\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(-1)^{(n-k)(k-1)}\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},$
$={\beta}\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},$ (158)
where we use $(n-1)k+(n-k)(k-1)=2nk-k(k+1)-(n-k)=n-k$ (mod 2). From the fact that eq. (109) holds for $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$, it is shown that $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ satisfies the rule (114) as follows,
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{i=1}^{n}(-1)^{n-i}(\nu_{n})_{l_{1}~{}\!\!\cdots~{}\!\!\hat{l}_{i}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{n+1}}=\sum_{i=1}^{n}(\nu_{n})_{l_{1}~{}\!\!\cdots~{}\!\!l_{i-1}~{}\!\!l_{n+1}~{}\!\!l_{i+1}~{}\!\!\cdots~{}\!\!l_{n}}$
$=(\nu_{n})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}+\cdots+(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}+(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}$
$=(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{n+1}}+(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}+\cdots+(\nu_{n})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}.$ (159)
Here, for completeness, we show that $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ given by eq. (108) have a cyclic property secretly. Using eq. (108), we obtain the relation:
$(\nu_{n}^{(0)})_{l_{p+1}~{}\!\!l_{p+2}~{}\!\!\cdots~{}\!\!l_{n}~{}\!\!l_{1}~{}\!\!\cdots~{}\!\!l_{p}}$ $={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{n-1}(E_{a_{r}})_{l_{r+p}~{}\!\!l_{p}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{p+1}~{}\!\!l_{p}}(E_{a_{2}})_{l_{p+2}~{}\!\!l_{p}}\cdots(E_{a_{n-p-1}})_{l_{n-1}~{}\!\!l_{p}}(E_{a_{n-p}})_{l_{n}~{}\!\!l_{p}}$
×$~{}~{}~{}~{}\times(E_{a_{n-p+1}})_{l_{1}~{}\!\!l_{p}}\cdots(E_{a_{n-1}})_{l_{p-1}~{}\!\!l_{p}}$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}((E_{a_{1}})_{l_{p+1}~{}\!\!l_{n}}+(E_{a_{1}})_{l_{n}~{}\!\!l_{p}})((E_{a_{2}})_{l_{p+2}~{}\!\!l_{n}}+(E_{a_{2}})_{l_{n}~{}\!\!l_{p}})\cdots$
×$~{}~{}~{}~{}\times((E_{a_{n-p-1}})_{l_{n-1}~{}\!\!l_{n}}+(E_{a_{n-p-1}})_{l_{n}~{}\!\!l_{p}})(E_{a_{n-p}})_{l_{n}~{}\!\!l_{p}}$
×$~{}~{}~{}~{}\times((E_{a_{n-p+1}})_{l_{1}~{}\!\!l_{n}}+(E_{a_{n-p+1}})_{l_{n}~{}\!\!l_{p}})\cdots((E_{a_{n-1}})_{l_{p-1}~{}\!\!l_{n}}+(E_{a_{n-1}})_{l_{n}~{}\!\!l_{p}})$
$={\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{p+1}~{}\!\!l_{n}}(E_{a_{2}})_{l_{p+2}~{}\!\!l_{n}}\cdots$
×$~{}~{}~{}~{}\times(E_{a_{n-p-1}})_{l_{n-1}~{}\!\!l_{n}}(-E_{a_{n-p}})_{l_{p}~{}\!\!l_{n}}(E_{a_{n-p+1}})_{l_{1}~{}\!\!l_{n}}\cdots(E_{a_{n-1}})_{l_{p-1}~{}\!\!l_{n}}$
$=-{\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{n-p+1}})_{l_{1}~{}\!\!l_{n}}\cdots(E_{a_{n-1}})_{l_{p-1}~{}\!\!l_{n}}$
×$~{}~{}~{}~{}\times(E_{a_{n-p}})_{l_{p}~{}\!\!l_{n}}(E_{a_{1}})_{l_{p+1}~{}\!\!l_{n}}(E_{a_{2}})_{l_{p+2}~{}\!\!l_{n}}\cdots(E_{a_{n-p-1}})_{l_{n-1}~{}\!\!l_{n}}$
$=(-1)^{(n-1)p}{\beta}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(E_{a_{1}})_{l_{1}~{}\!\!l_{n}}\cdots(E_{a_{p-1}})_{l_{p-1}~{}\!\!l_{n}}$
×$~{}~{}~{}~{}\times(E_{a_{p}})_{l_{p}~{}\!\!l_{n}}(E_{a_{p+1}})_{l_{p+1}~{}\!\!l_{n}}(E_{a_{p+2}})_{l_{p+2}~{}\!\!l_{n}}\cdots(E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n}}$
$=(-1)^{(n-1)p}(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (160)
where the factor $-(-1)^{(n-1)p}$ comes from the $(n-2)(n-p-1)+(p-1)(=(n-1)p+1({\rm mod}~{}2))$ exchange of the indices on $E_{a_{r}}$ from $(a_{n-p+1},\cdots,a_{n-1},a_{n-p},a_{1},a_{2},\cdots,a_{n-p-1})$ to $(a_{1},\cdots,a_{n-p-2},a_{n-p-1},a_{n-p},a_{n-p+1},\cdots,a_{n-1})$.
Finally, let us show that a variable given by
$A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(0)e^{2\pi i(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}t}$ (161)
with the frequency:
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ $=\frac{\gamma}{h}(-1)^{\frac{(n-2)(n-3)}{2}+1}\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}$
×$~{}~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (162)
satisfies the generalized Heisenberg equation:
$\frac{d}{dt}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (163)
where $(H_{a_{k}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ ($k=1,2\cdots,n-1$) are the Hamiltonians with the normal forms whose non-vanishing components are given by
$(H_{a_{k}})_{l_{1}~{}\!\!\cdots~{}\!\!l_{i}~{}\!\!\cdots~{}\!\!l_{j-1}~{}\!\!l_{i}~{}\!\!l_{j+1}~{}\!\!\cdots~{}\!\!l_{n}}=\sum_{u(\neq i,j)}(h_{a_{k}})_{l_{u}~{}\!\!l_{i}},$ (164)
with constants $(h_{a_{k}})_{l_{u}~{}\!\!l_{i}}$. In eq. (164), $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, and $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different. We notice that eq. (162) agrees with eq. (158) by identifying ${\beta}$, $(E_{a_{r}})_{l_{r}~{}\!l_{k}}$ and $(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!l_{k}}$ as
${\beta}=\frac{\gamma}{h}(-1)^{\frac{(n-2)(n-3)}{2}+1},~{}~{}$ (165)
$(E_{a_{r}})_{l_{r}~{}\!l_{k}}=(h_{a_{r}})_{l_{r}~{}\!l_{k}}~{}~{}(r=1,\cdots,k-1),~{}~{}$ (166)
$(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!l_{k}}=(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!l_{k}}~{}~{}(r^{\prime}=k,\cdots,n-1),$ (167)
respectively.
First, exchanging the Hamiltonians, we obtain the relation:
$\left[A(t),H_{1},H_{2},\cdots,H_{n-2},H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ (168)
$~{}~{}~{}=(-1)^{\frac{(n-2)(n-3)}{2}}\left[A(t),H_{n-2},\cdots,H_{2},H_{1},H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}.$
The first term $(A(t)H_{n-2}H_{n-3}\cdots H_{2}H_{1}H_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ in the right-hand side of eq. (168) is calculated as
$(A(t)H_{n-2}H_{n-3}\cdots H_{2}H_{1}H_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}=\sum_{l_{n+1}}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}(H_{n-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}(H_{n-3})_{l_{1}~{}\!\!\cdots~{}\!\!l_{n-3}~{}\!\!l_{n+1}~{}\!\!l_{n-1}~{}\!\!l_{n}}$
××$~{}~{}~{}~{}~{}~{}\times\cdots\times(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{n+1}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-1})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}=A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n}~{}\!\!l_{n}}(H_{n-3})_{l_{1}~{}\!\!\cdots~{}\!\!l_{n-3}~{}\!\!l_{n}~{}\!\!l_{n-1}~{}\!\!l_{n}}$
××$~{}~{}~{}~{}~{}~{}\times\cdots\times(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{n}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{n}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-1})_{l_{n}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}=A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\left\{\prod_{r=1}^{n-1}(h_{r})_{l_{r}~{}\!\!l_{n}}+\sum_{k=2}^{n-2}\prod_{r=1}^{n-k-1}(h_{r})_{l_{r+k}~{}\!\!l_{n}}\prod_{r^{\prime}=n-k}^{n-1}(h_{r^{\prime}})_{l_{r^{\prime}+k+1}~{}\!\!l_{n}}+\cdots\right\},$ (169)
where we use the relation:
$(H_{k})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k}~{}\!\!l_{n}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{u(\neq k+1,n)}(h_{k})_{l_{u}~{}\!\!l_{n}}$
$=(h_{k})_{l_{1}~{}\!\!l_{n}}+\cdots+(h_{k})_{l_{k}~{}\!\!l_{n}}+(h_{k})_{l_{k+2}~{}\!\!l_{n}}+\cdots+(h_{k})_{l_{n-1}~{}\!\!l_{n}}$ (170)
for $k=1,2,\cdots,n-2$ and the relation:
$(H_{n-1})_{l_{n}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{u(\neq 1,n)}(h_{k})_{l_{u}~{}\!\!l_{n}},$
$=(h_{n-1})_{l_{2}~{}\!\!l_{n}}+(h_{n-1})_{l_{3}~{}\!\!l_{n}}+\cdots+(h_{n-1})_{l_{n-1}~{}\!\!l_{n}}.$ (171)
In eqs. (170) and (171), $(l_{1},\cdots,l_{k},l_{k+2},\cdots l_{n})$ and $(l_{2},l_{3},\cdots,l_{n})$ are all different, respectively. The ellipsis in eq. (169) stands for terms including both $(h_{a_{k}})_{l_{r}~{}\!\!l_{n}}$ and $(h_{a_{k^{\prime}}})_{l_{r}~{}\!\!l_{n}}$ ($a_{k}\neq a_{k^{\prime}}$) and all of them are eliminated by the addition of the terms which come from the exchange of $H_{a_{k}}$ and $H_{a_{k^{\prime}}}$.
Actually, by adding all the terms with the Hamiltonians interchanged, we obtain the relation:
$\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(A(t)H_{a_{n-2}}H_{a_{n-3}}\cdots H_{a_{2}}H_{a_{1}}H_{a_{n-1}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}=\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\left\{\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{n}}+\sum_{k=2}^{n-2}\prod_{r=1}^{n-k-1}(h_{a_{r}})_{l_{r+k}~{}\!\!l_{n}}\prod_{r^{\prime}=n-k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+k+1}~{}\!\!l_{n}}\right\}$
$~{}~{}~{}=\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\left\{\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{n}}+\sum_{k=2}^{n-2}(-1)^{k(n-2)}\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{n}}\right\}$
$~{}~{}~{}=\gamma\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{n}}~{}\!A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (172)
where $\gamma=1$ for an odd number $n(\geq 3)$ and $\gamma=n-2$ for an even number $n(\geq 4)$, and we use that ${\prod_{r=1}^{n-k-1}(h_{a_{r}})_{l_{r+k}~{}\!\!l_{n}}\prod_{r^{\prime}=n-k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+k+1}~{}\!\!l_{n}}}$ are replaced by ${(-1)^{k(n-2)}\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{n}}}$ after exchanging $(a_{1},a_{2},\cdots,a_{n-1})$ for $(a_{k+1},a_{k+2},\cdots,a_{k})$.
In the same way, the term $(-1)^{n-2-k}(H_{n-2}H_{n-3}\cdots H_{k-1}A(t)H_{k-2}\cdots H_{2}H_{1}H_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ is calculated as
$(-1)^{n-2-k}(H_{n-2}H_{n-3}\cdots H_{k-1}A(t)H_{k-2}\cdots H_{2}H_{1}H_{n-1})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}~{}=(-1)^{n-k}\sum_{l_{n+1}}(H_{n-2})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n+1}}(H_{n-3})_{l_{1}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{n+1}~{}\!\!l_{n}}\cdots(H_{k-1})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k}~{}\!\!l_{n+1}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}}$
×$~{}~{}~{}~{}~{}~{}~{}~{}\times A(t)_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-1}~{}\!\!l_{n+1}~{}\!\!l_{k+1}~{}\!\!\cdots~{}\!\!l_{n}}(H_{k-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-2}~{}\!\!l_{n+1}~{}\!\!l_{k}~{}\!\!\cdots~{}\!\!l_{n}}\cdots(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{n+1}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}$
×$~{}~{}~{}~{}~{}~{}~{}~{}\times(H_{n-1})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}~{}=(-1)^{n-k}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{k-1}~{}\!\!l_{k}~{}\!\!l_{k+1}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-2})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-1}~{}\!\!l_{k}}(H_{n-3})_{l_{1}~{}\!\!\cdots~{}\!\!l_{n-2}~{}\!\!l_{k}~{}\!\!l_{n}}\cdots(H_{k-1})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k}~{}\!\!l_{k}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}}$
×$~{}~{}~{}~{}~{}~{}~{}~{}\times(H_{k-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-2}~{}\!\!l_{k}~{}\!\!l_{k}~{}\!\!\cdots~{}\!\!l_{n}}\cdots(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{k}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{k}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-1})_{l_{k}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}~{}=(-1)^{n-k}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\left\{\prod_{r=1}^{k-1}(h_{r})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{r^{\prime}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{m=2}^{k-3}\prod_{r=1}^{k-1-m}(h_{r})_{l_{r+m}~{}\!\!l_{k}}\prod_{r^{\prime}=k-m}^{n-1}(h_{r^{\prime}})_{l_{r^{\prime}+m+1}~{}\!\!l_{k}}+(h_{1})_{l_{k-1}~{}\!\!l_{k}}\prod_{r=2}^{n-1}(h_{r})_{l_{r+k-1}~{}\!\!l_{k}}+\prod_{r=1}^{n-1}(h_{r})_{l_{r+k}~{}\!\!l_{k}}$
$~{}~{}~{}~{}~{}~{}~{}~{}\left.+\sum_{m={k+1}}^{n-1}\prod_{r=1}^{k-1-m}(h_{r})_{l_{r+m}~{}\!\!l_{k}}\prod_{r^{\prime}=k-m}^{n-1}(h_{r^{\prime}})_{l_{r^{\prime}+m+1}~{}\!\!l_{k}}+\cdots\right\},$ (173)
where we use the relation:
$(H_{j})_{l_{1}~{}\!\!\cdots~{}\!\!l_{j}~{}\!\!l_{k}~{}\!\!l_{j+2}~{}\!\!\cdots~{}\!\!l_{k-1}~{}\!\!l_{k}~{}\!\!l_{k+1}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{u(\neq j+1,k)}(h_{j})_{l_{u}~{}\!\!l_{k}}$
$=(h_{j})_{l_{1}~{}\!\!l_{k}}+\cdots+(h_{j})_{l_{j}~{}\!\!l_{k}}+(h_{j})_{l_{j+2}~{}\!\!l_{k}}$
$~{}~{}~{}~{}+\cdots+(h_{j})_{l_{k-1}~{}\!\!l_{k}}+(h_{j})_{l_{k+1}~{}\!\!l_{k}}+\cdots+(h_{j})_{l_{n}~{}\!\!l_{k}}$ (174)
for $31\leq j\leq k-$ 3,
$(H_{k-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k}~{}\!\!l_{k}~{}\!\!l_{k+1}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{u(\neq k-1,k)}(h_{k-2})_{l_{u}~{}\!\!l_{k}}$
$=(h_{k-2})_{l_{1}~{}\!\!l_{k}}+\cdots+(h_{k-2})_{l_{k-2}~{}\!\!l_{k}}+(h_{k-2})_{l_{k+1}~{}\!\!l_{k}}+\cdots+(h_{k-2})_{l_{n}~{}\!\!l_{k}},$ (175)
$(H_{k-1})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k}~{}\!\!l_{k}~{}\!\!l_{k+2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{u(\neq k,k+1)}(h_{k-1})_{l_{u}~{}\!\!l_{k}}$
$=(h_{k-1})_{l_{1}~{}\!\!l_{k}}+\cdots+(h_{k-1})_{l_{k-1}~{}\!\!l_{k}}+(h_{k-1})_{l_{k+2}~{}\!\!l_{k}}+\cdots+(h_{k-1})_{l_{n}~{}\!\!l_{k}}$ (176)
and
$(H_{j})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-1}~{}\!\!l_{k}~{}\!\!l_{k+1}~{}\!\!\cdots~{}\!\!l_{j+1}~{}\!\!l_{k}~{}\!\!l_{j+3}~{}\!\!\cdots~{}\!\!l_{n}}$ $=\sum_{u(\neq k,j+2)}(h_{j})_{l_{u}~{}\!\!l_{k}}$
$=(h_{j})_{l_{1}~{}\!\!l_{k}}+\cdots+(h_{j})_{l_{k-1}~{}\!\!l_{k}}+(h_{j})_{l_{k+1}~{}\!\!l_{k}}$
$~{}~{}~{}~{}+\cdots+(h_{j})_{l_{j+1}~{}\!\!l_{k}}+(h_{j})_{l_{j+3}~{}\!\!l_{k}}+\cdots+(h_{j})_{l_{n}~{}\!\!l_{k}}$ (177)
for $k\leq j\leq n-1$. In eqs. (174), (175), (176) and (177), the $n-1$ indices $(l_{1},\cdots,l_{j},l_{j+2},\cdots l_{n})$, $(l_{1},\cdots,l_{k-2},l_{k},\cdots l_{n})$, $(l_{1},\cdots,l_{k},l_{k+2},\cdots l_{n})$, and $(l_{1},\cdots,l_{j+1},l_{j+3},\cdots l_{n})$ are all different, respectively. The ellipsis in eq. (173) stands for terms including both $(h_{a_{k}})_{l_{r}~{}\!\!l_{n}}$ and $(h_{a_{k^{\prime}}})_{l_{r}~{}\!\!l_{n}}$ ($a_{k}\neq a_{k^{\prime}}$) and all of them are eliminated by adding the terms which stem from the exchange of $H_{a_{k}}$ and $H_{a_{k^{\prime}}}$.
By adding all the terms with the Hamiltonians interchanged, we obtain the relation:
$(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}(H_{a_{n-2}}H_{a_{n-3}}\cdots H_{a_{k-1}}A(t)H_{a_{k-2}}\cdots H_{a_{2}}H_{a_{1}}H_{a_{n-1}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$~{}~{}~{}=(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\left\{\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}\right.$
$~{}~{}~{}~{}~{}~{}~{}+\sum_{m=2}^{k-3}\prod_{r=1}^{k-1-m}(h_{a_{r}})_{l_{r+m}~{}\!\!l_{k}}\prod_{r^{\prime}=k-m}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+m+1}~{}\!\!l_{k}}+(h_{a_{1}})_{l_{k-1}~{}\!\!l_{k}}\prod_{r=2}^{n-1}(h_{a_{r}})_{l_{r+k-1}~{}\!\!l_{k}}+\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r+k}~{}\!\!l_{k}}$
$~{}~{}~{}~{}~{}~{}~{}\left.+\sum_{m={k+1}}^{n-1}\prod_{r=1}^{k-1-m}(h_{a_{r}})_{l_{r+m-1}~{}\!\!l_{k}}\prod_{r^{\prime}=k-m}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+m+1}~{}\!\!l_{k}}+\cdots\right\}$
$~{}~{}~{}=(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\left\{\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}\right.$
$~{}~{}~{}~{}~{}~{}~{}\left.+\sum_{m=2}^{n-2}(-1)^{m(n-2)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}\right\}$
$~{}~{}~{}=\gamma(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}~{}\!A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (178)
where $\gamma=1$ for an odd number $n(\geq 3)$ and $\gamma=n-2$ for an even number $n(\geq 4)$, and we use that the factors such as
$\sum_{m=2}^{k-3}\prod_{r=1}^{k-1-m}(h_{a_{r}})_{l_{r+m}~{}\!\!l_{k}}\prod_{r^{\prime}=k-m}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+m+1}~{}\!\!l_{k}},~{}~{}(h_{a_{1}})_{l_{k-1}~{}\!\!l_{k}}\prod_{r=2}^{n-1}(h_{a_{r}})_{l_{r+k-1}~{}\!\!l_{k}},$
$\prod_{r=1}^{n-1}(h_{a_{r}})_{l_{r+k}~{}\!\!l_{k}},~{}~{}\sum_{m={k+1}}^{n-1}\prod_{r=1}^{k-1-m}(h_{a_{r}})_{l_{r+m-1}~{}\!\!l_{k}}\prod_{r^{\prime}=k-m}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+m+1}~{}\!\!l_{k}}$ (179)
are replaced by
$\sum_{m=2}^{k-3}(-1)^{m(n-2)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},~{}~{}(-1)^{(k-2)(n-2)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},$
$(-1)^{(k-1)(n-2)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},~{}~{}\sum_{m={k+1}}^{n-1}(-1)^{(m-1)(n-2)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},$ (180)
respectively, after exchanging $(a_{1},a_{2},\cdots,a_{n-1})$ for suitable ones. We notice that eq. (172) corresponds to the case with $k=n$ in eq. (178).
The left-hand side of eq. (163) becomes as
$\frac{d}{dt}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=2\pi i(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t).$ (181)
On the other hand, the right-hand side of eq. (163) becomes as
$\frac{1}{i\hbar}\left[A(t),H_{1},H_{2},\cdots,H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ $=-\frac{2\pi i}{h}(-1)^{\frac{(n-2)(n-3)}{2}}\left[A(t),H_{n-2},\cdots,H_{2},H_{1},H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
$=\frac{2\pi i}{h}(-1)^{\frac{(n-2)(n-3)}{2}+1}\gamma\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}$
×$~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}~{}\!A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (182)
where we use eqs. (168) and (178). From eqs. (181) and (182), eq. (162) is obtained and hence it is shown that $A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ given eqs. (161) and (162) satisfies eq. (163).
Incidentally, we point out that, if the generalized Heisenberg equation is defined by
$\frac{d}{dt}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=\frac{1}{i\hbar}\left[A(t),H_{n-2},\cdots,H_{2},H_{1},H_{n-1}\right]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (183)
the formula of $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ takes the form:
$(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}=-\frac{\gamma}{h}\sum_{k=1}^{n}(-1)^{n-k}\sum_{(a_{1},a_{2},\cdots,a_{n-1})}{\rm sgn(P)}\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (184)
due to the disappearance of the factor $(-1)^{\frac{(n-2)(n-3)}{2}}$, as seen from eqs. (162), (163) and (168).